### v1

# ObjectClear: Complete Object Removal via Object-Effect Attention

Abstract
Object removal requires eliminating not only the target object but also its effects, such as shadows and reflections. However, diffusion-based inpainting methods often produce artifacts, hallucinate content, alter background, and struggle to remove object effects accurately. To address this challenge, we introduce a new dataset for OBject-Effect Removal, named OBER, which provides paired images with and without object effects, along with precise masks for both objects and their associated visual artifacts. The dataset comprises high-quality captured and simulated data, covering diverse object categories and complex multi-object scenes. Building on OBER, we propose a novel framework, ObjectClear, which incorporates an object-effect attention mechanism to guide the model toward the foreground removal regions by learning attention masks, effectively decoupling foreground removal from background reconstruction. Furthermore, the predicted attention map enables an attention-guided fusion strategy during inference, greatly preserving background details. Extensive experiments demonstrate that ObjectClear outperforms existing methods, achieving improved object-effect removal quality and background fidelity, especially in complex scenarios.
footnotetext: Corresponding author.
1 Introduction
Figure 1: Object Removal Comparison. Given an object mask, prior methods often leave residual artifacts or hallucinate undesirable content, change background, and typically fail to remove associated effects such as shadows and reflections. In contrast, our ObjectClear precisely eliminates both the object and its associated effects, achieving seamless object removal results even in challenging cases.
Recent advances in generative models [27, 24, 21], such as GPT-4o, have shown strong capabilities in image editing and are widely adopted in real-world applications. Among these tasks, object removal has become a key task, allowing users to erase unwanted content from images. However, seamlessly removing both the object and its effects (e.g., shadows and reflections) while preserving the background remains a challenging problem.
While diffusion-based methods [1, 43, 12, 15, 31, 7, 10, 5] have advanced object removal, they still struggle with high-fidelity results. As shown in Fig. 1, existing methods often leave residual artifacts, hallucinate unwanted content, alter backgrounds details, and fail to remove object effects. This is partly because the input mask fully occludes the target object, preventing models from perceiving its surrounding context and associated effects. In addition, the lack of large-scale and public datasets for object-effect removal also limits further progress in the field.
Existing object removal datasets can be categorized into simulated and camera-captured data. (1) Simulated data. These datasets are often constructed by copy-pasting objects [11, 14] or using pretrained inpainting models to generate pseudo ground truth [32]. While this allows low-cost generation of large-scale data, such datasets typically lack object effects such as shadows and reflections, causing models trained on them to struggle with effect removal. (2) Camera-captured data. Some works [29, 33] leverage fixed-viewpoint videos to extract paired frames, but this constrains the foreground types to moving objects. Others [34, 39, 36] capture image pairs before and after object removal, but these datasets are often not publicly available and small in scale due to the high costs.
To bridge this gap, we introduce OBER, a hybrid dataset for OBject-Effect Removal, which combines both camera-captured and simulated data with diverse foreground objects, background scenes, and object effects. For the camera-captured data, we annotate precise object and object-effect masks, which serve as critical supervision during training. For the simulated data, we leverage these masks to compute accurate alpha maps, enabling realistic alpha blending of RGBA object layers with effects and high-quality backgrounds. We further extend the simulation pipeline to include multi-object scenarios, enhancing model robustness in challenging cases involving occlusion and object interactions, as shown in the last two rows of Fig. 5. By combining the realism of captured data with the scalability of simulation, OBER offers a high-quality, large-scale, and diverse dataset with a total of 12,715 training samples. In addition, we introduce two new benchmarks: OBER-Test (w/ ground truth) and OBER-Wild (w/o ground truth), to support future research.
Built upon OBER, we propose a novel framework, ObjectClear, which introduces an Object-Effect Attention (OEA) mechanism that adaptively focuses on the foreground removal region, including both the object and its associated effects, by supervising cross-attention maps with object-effect masks. The OEA module effectively decouples foreground removal from background reconstruction, enabling more precise and complete object elimination. In addition, the predicted attention map supports an Attention-Guided Fusion (AGF) strategy during inference, helping to preserve background details, as shown in Fig. 7. This design enables the model to localize editable regions adaptively and offers a potential solution to the long-standing fidelity issues in diffusion-based image editing.
Our contributions are summarized as follows:
  • We propose OBER, a high-quality and large-scale hybrid dataset for object removal, featuring diverse objects, fine-grained annotations of object-effect masks, and complex multi-object scenarios across both simulated and camera-captured settings.
  • We introduce a novel framework, ObjectClear, which incorporates an Object-Effect Attention mechanism that adaptively focuses on foreground removal regions and enables an Attention-Guided Fusion strategy, improving object removal quality and background fidelity.
  • We present two new benchmarks, OBER-Test and OBER-Wild, providing a foundation for future research. Our method achieves superior performance on all benchmarks, outperforming existing approaches in terms of both quantitative metrics and visual quality.
2 Related Work
Image Inpainting. Image inpainting is a long-standing visual editing task that aims to seamlessly reconstruct pixels within a given mask. Early approaches predominantly adopt generative adversarial networks (GANs) [8], but often suffer from limited realism and diversity [17, 22, 26, 40]. With the rapid advances in diffusion models [28, 30, 9, 24], many methods [2, 20, 21, 43, 12, 38] have begun leveraging their strong generative priors to synthesize high-fidelity content, achieving state-of-the-art results in image inpainting. In this work, we adapt the SDXL-Inpainting model [1] for photorealistic completion. However, despite their strong generative capabilities, existing inpainting models often lack awareness of object-induced effects (e.g., shadows and reflections), leading to incomplete or visually inconsistent object removal results.
Object Removal. Object removal is a specialized form of image inpainting that requires accounting for object effects to ensure complete and realistic removal, which diffusion-based models [34, 11, 19, 39, 15, 4, 43, 7, 5, 36, 10, 31, 33] currently dominate. A common strategy involves curating high-quality triplet datasets. For instance, ObjectDrop [34] builds a real-world dataset by having photographers capture the same scene before and after the removal of a single object. However, its limited scale and lack of public availability hinder broader adoption. To address data scalability, methods like SmartEraser [11] and Erase Diffusion [19] rely on synthetic datasets using segmentation or matting techniques to extract foreground objects, but these typically lack annotations of object effects, limiting the models’ ability to remove shadows or reflections. To enhance realism, methods such as LayerDecomp [36] and OmniPaint [39] create costly camera-captured data. OmniPaint [39] takes one step further to annotate unlabeled images by a model trained on small-scale real data, while RORem [15] involves human annotators to ensure the quality of annotated data. In parallel, works like RORD [29] and OmniEraser [33] scale data generation by mining realistic video frames with fixed viewpoints, selectively pairing frames with and without target objects while preserving natural object effects. To eliminate reliance on curated datasets altogether, other methods opt for test-time optimization [5, 31, 10] or providing plug-and-play solutions [7]. However, all these models rely on implicit learning of object effects without explicitly modeling effect maps, making it difficult to ensure that no unnecessary background changes occur.
3 Methodology
To achieve accurate removal of target objects along with their associated visual effects, such as shadow and reflection, we propose a comprehensive framework consisting of a data curation pipeline (OBER Dataset) and a dedicated object removal model (ObjectClear). These two components are detailed in Sec. 3.1 and Sec. 3.2, respectively.
3.1 OBER Dataset
ObjectDrop [34] shows that even small-scale datasets, when captured with careful camera control over scene consistency, can significantly enhance model generalizability in object removal tasks. However, due to the lack of public access to such datasets and the high cost of real-world data collection, subsequent works often face limitations in dataset scale. To address this challenge, we introduce the OBER dataset (OBject-Effect Removal), designed to balance data realism and scalability. As illustrated in Fig. 2, OBER is a hybrid dataset consisting of two parts: (1) a small set of camera-captured images adhering to physical realism, and (2) a larger set of simulated images generated by compositing foreground objects, extracted from the real data in (1), onto diverse background scenes.
Figure 2: Dataset Construction Pipeline of OBER. The dataset combines both camera-captured and simulated data, featuring diverse foreground objects and background scenes. It provides rich annotations, including object masks, object-effect masks, transparent RGBA object layers, and complex multi-object scenarios for training and evaluation.
Camera-Captured Data. (1) Capture Paired Images. Following the approach of ObjectDrop [34], we use fixed cameras to construct a counterfactual dataset by capturing each scene before and after the removal of a single object, while keeping all the other factors unchanged. For each pair, the image with the object is used as the input $I_{in}$, and the image without the object serves as the ground truth $I_{gt}$. In total, we collected 2,875 such counterfactual pairs with 2,715 for training and 160 for testing. Our dataset encompasses a wide variety of everyday objects commonly found in indoor (e.g., benches, cups, dolls) and outdoor scenes (e.g., pedestrians, vehicles). These image pairs also preserve visual effects such as shadows and reflections. To reduce potential pixel-level misalignment, we downsample and crop all images to a fixed training resolution (i.e., 512 $\times$ 512). (2) Segment Object Mask $M_{o}$. To obtain the object mask $M_{o}$ as network input alongside the input image $I_{in}$, we apply off-the-shelf detection and segmentation models, such as DINO [41] and SAM [13] to $I_{in}$. (3) Segment Object-Effect Mask $M_{o}\cup M_{e}$. Different from previous methods, we propose to explicitly model the object effects. Therefore, we introduce an object-effect mask $M_{fg}$ that covers both the object $M_{o}$ and its associated visual effects $M_{e}$ in our dataset. Unlike the coarse object-effect masks provided in the RORD dataset [29], we construct $M_{fg}$ in an efficient and accurate way by computing the pixel-wise difference between $I_{in}$ and $I_{gt}$. Pixels with differences above a predefined threshold are regarded as part of the object-effect mask. This object-effect mask provides crucial supervision during training, allowing the network to adaptively learn to focus on the target removal regions, including both the object itself and its effects.
Simulated Data. With the high-quality real data collected, we further scale up the dataset with a carefully designed simulation pipeline. (1) Collect Background Images. We begin by downloading high-quality background images from the Internet, which offer a diverse range of background scenes. To select backgrounds with flat regions suitable for object placement, we first apply Mask2Former [6] to segment flat areas corresponding to semantic classes such as “road, sidewalk, grass, floor”, as illustrated in Fig. 2(b). We then refine the selection by computing the gradient of the depth map generated by Depth Anything V2 [37], filtering out regions with significant depth variation to ensure that the inserted objects are placed on visually flat surfaces. (2) Collect Foreground Objects with Effects. Based on the camera-captured paired data $(I_{in},I_{gt})$, along with the object mask $M_{o}$ and effect mask $M_{e}$, we compute the alpha map of the foreground object with effects $I_{oe}$ using Eq. 1, where $\varepsilon$ is a small constant added to avoid division by zero. For subsequent compositing, we manually categorize the foreground objects into eight groups based on their shadow direction (Fig. 2(b)).
$\alpha(p)=\begin{cases}0,&\text{if pixel }p\in\text{background area $\overline{M_{o}}\cap\overline{M_{e}}$}\\ 1,&\text{if pixel }p\in\text{object area $M_{o}$}\\ (I_{\text{gt}}-I_{\text{in}})/(I_{\text{gt}}+\varepsilon),&\text{if pixel }p\in\text{effect area $M_{e}$}\\ \end{cases}$ (1)
(3) Blend Objects with Backgrounds. We randomly sample a background image $I_{bg}$ and a foreground object image with effects $I_{oe}$, and synthesize a composite image using alpha blending: $I_{comp}=(1-\alpha)\cdot I_{bg}+\alpha\cdot I_{oe}$. Beyond single-object cases, we also synthesize multi-object data by compositing multiple foreground objects with the same lighting directions, thereby covering scenarios involving object occlusions, as illustrated in Fig. 2(b). In total, we generate 10,000 composite images, significantly enriching the diversity and scalability of our dataset. In particular, the simulation of multi-object compositions leads to notable improvements in object removal robustness and background preservation, as discussed and compared in Table 2.
3.2 ObjectClear
Figure 3: The Framework of ObjectClear. Given an input image and a target object mask, ObjectClear employs an Object-Effect Attention mechanism to guide the model toward foreground removal regions by learning attention masks. The predicted mask further enables an Attention-Guided Fusion strategy during inference, which substantially preserves background details.
Our ObjectClear is built upon SDXL-Inpainting [1]. While SDXL-Inpainting uses a noise map $z_{t}$, a masked image $I_{m}$, a corresponding mask $M_{o}$, and a text prompt $c$ as inputs, we feed the original image $I_{in}$ instead of $I_{m}$ into the model, and our inputs are expressed as a tuple $<z_{t},I_{in},M_{o},c>$. This design encourages the model to better attend to the effect of the target object by leveraging its visual features. Moreover, it facilitates more effective utilization of background information behind the object in cases involving moving transparent objects, such as glass cups. Apart from the input design, we further introduce an Object-Effect Attention Mechanism and an Attention-Guided Fusion Strategy to explicitly attend object-effect regions and enhance background preservation. Detailed descriptions of these components are provided below.
Object-Effect Attention. To enable the model to better attend to both the target object region and its associated visual effect regions, we follow the approach of Li et al. [16] by stacking text and visual object information as guidance for the cross-attention layers in the base model. Specifically, the text prompt is expressed as “remove the instance of”, and the visual object is obtained by applying a dot product between the input image $I_{in}$ and $M_{o}$, as illustrated in Fig. 3(a). These two modalities are then encoded into text embeddings and visual embeddings using the CLIP [25] text and vision encoders, respectively. To unify their representation spaces, the visual embeddings are further projected into the same dimensional space as the text embeddings using a lightweight trainable MLP composed of two linear layers. The resulting text embeddings and projected object embeddings are then stacked and used as guidance in the cross-attention blocks of the base model. To encourage the model to focus more accurately on the object and its effect regions, we introduce a mask loss $\mathcal{L}_{mask}$. Concretely, we extract the cross-attention maps corresponding to the visual embedding tokens and denote them as $\mathbf{A}$, which we supervise with the annotated foreground object-effect masks $M_{fg}$ from our OBER dataset. Following [35, 23], $\mathcal{L}_{\text{mask}}$ is designed to minimize the attention values in the background regions while maximizing those in the foreground. This objective can be formulated as:
$\mathcal{L}_{mask}=\mathrm{mean}({\mathbf{A}}[1-M_{fg}])-\mathrm{mean}({\mathbf{A}}[M_{fg}])\text{,}$ (2)
where []delimited-[][$\cdot$][ ⋅ ] denotes indexing operator, and $\mathbf{A}[M_{fg}]$ refers to the attention values within the foreground regions specified by the mask $M_{fg}$.
Attention-Guided Fusion. Interestingly, we observe that applying Object-Effect Attention not only improves ObjectClear’s precision in object removal, but also produces attention maps that accurately capture both the object and its effects. Based on this observation, we propose Attention-Guided Fusion strategy that utilizes attention maps during inference to seamlessly combine the generated result with the original image. Specifically, we extract the first-layer cross-attention map corresponding to the object embedding during inference. This attention map is then upsampled to match the spatial resolution of the original image, forming a soft estimate of the object-effect region. To enhance the spatial smoothness and avoid hard edges, we apply a Gaussian blur to the upsampled attention map, producing a softened object-effect mask. This blurred mask is then used to blend the generated image with the original input via a copy-and-paste operation, where pixels within the object-effect region are taken from the generated result, and the rest are preserved from the original image. This strategy significantly reduces undesired background changes caused by the diffusion process and VAE reconstruction, while preserving high-fidelity object removal. Unlike BrushNet [12], which relies on user-provided object masks to guide blending, our approach adaptively produces the object-effect mask from our Object-Affect Attention.
4 Experiments
Table 1: Quantitative Comparisons with State-of-the-Art Methods. The best and second performances are marked in red and orange, respectively. Although ObjectClear takes only the object mask as input, it outperforms previous methods, even when those methods are provided with masks that cover both the object and its associated effect regions.
Mask Datasets RORD-Val OBER-Test OBER-Wild
[…]
Object-Effect Mask SDXL-INP [1] 19.39 22.98 0.2432 0.1024 24.07 26.93 0.1296 0.0579 0.8356
PowerPaint [43] 19.87 21.98 0.2303 0.0776 26.20 27.41 0.1243 0.0409 0.8878
BrushNet [12] 16.82 18.87 0.3434 0.1692 20.96 23.69 0.2052 0.1180 0.6415
DesignEdit [10] 20.69 22.49 0.2946 0.1200 26.59 27.63 0.1777 0.0629 0.9150
CLIPAway [7] 18.87 20.78 0.3328 0.0969 25.38 26.28 0.1039 0.0349 0.8646
FreeCompose [5] 19.67 21.72 0.3316 0.1110 23.39 25.15 0.1305 0.0629 0.8976
Attentive Eraser [31] 20.33 21.98 0.2545 0.1015 27.42 29.27 0.1114 0.0249 0.9095
RORem [15] 21.61 23.11 0.3224 0.0767 27.23 27.95 0.1042 0.0234 0.9157
Object Mask SDXL-INP [1] 20.23 24.83 0.2042 0.0868 22.42 25.77 0.1428 0.0771 0.8044
PowerPaint [43] 21.46 24.62 0.1801 0.0648 22.76 24.67 0.1544 0.0729 0.8538
BrushNet [12] 18.06 23.44 0.2757 0.1821 21.19 24.38 0.1822 0.1123 0.7296
DesignEdit [7] 22.09 24.26 0.2501 0.1021 24.63 25.48 0.1870 0.0788 0.9002
CLIPAway [10] 20.58 23.21 0.2770 0.0785 22.32 24.05 0.1357 0.0765 0.8646
FreeCompose [5] 20.39 22.91 0.3015 0.0897 22.77 24.46 0.1393 0.0690 0.8688
Attentive Eraser [31] 22.17 24.59 0.1883 0.0643 25.70 27.08 0.1201 0.0437 0.9008
RORem [15] 22.49 24.10 0.2943 0.0634 24.51 25.28 0.1288 0.0460 0.9025
ObjectClear (Ours) 26.40 30.75 0.0802 0.0350 32.98 35.14 0.0352 0.0099 0.9172
Implementation. Our proposed method is built on SDXL-Inpainting [1] and fine-tuned with our OBER dataset with an input resolution of $512\times 512$. Training is conducted with a total batch size of $32$ on 8 A100 GPUs for 100k iterations and a learning rate of 1e-5. All experiment results are attained with a guidance scale of $7.5$ and $20$ denoising steps.
Evaluation Data. We evaluate the object removal performance of our method, ObjectClear, against existing approaches on three test datasets: (1) RORD-Val: RORD [29] is a widely used object removal dataset. Each sample contains an image pair (images with and without the target object) and a coarse mask covering both the target object and its associated effect. To avoid duplicated scenarios, we randomly select one image per scene, resulting in a RORD-Val with 343 samples. Since most of the object removal methods, including ObjectClear, require only the object mask as input, we augment RORD-Val with accurate object masks manually to enable a more comprehensive evaluation. (2) OBER-Test: As described in Sec. 3.1, we split our collected OBER dataset into a training set and a test set. The test set contains 160 samples, each comprising image pairs along with precise masks for both the object and its effect. (3) OBER-Wild: To further assess the performance of different methods on out-of-distribution scenarios in the wild, we collect 300 high-quality images featuring objects with associated effects (shadows or reflections) from the Internet. We annotate the object masks using DINO [41] and SAM [13], and manually annotate the effect masks. Note that the OBER-Wild set does not include the removal ground truth.
Figure 4: Object Removal on OBER-Test and RORD-Val. Our ObjectClear effectively removes both the masked objects and their associated effects.
Figure 5: Object Removal on OBER-Wild. Our ObjectClear not only effectively removes single objects with their shadows and reflections (top five samples), but also accurately removes the target object when multiple mutually occluding objects exist (bottom two samples).
Evaluation Metrics. For RORD-Val and OBER-Test, where ground truth is available, we evaluate performance using fidelity metrics: PSNR and PSNR-BG (computed only on background regions), as well as two perceptual metrics: LPIPS [42] and CLIP [25] (feature distance). For OBER-Wild, which lacks ground truth, we employ ReMOVE [3] to measure the consistency between the object removal region and the surrounding background.
Figure 6: Object-Effect Attention Map. ObjectClear achieves a relatively high recall in the object-effect attention map, with recall values increasing as the denoising step progresses. The attention map obtained in the final step effectively covers both the target object and its associated effects, including the object’s shadow (red box) and its reflection in the mirror (green box).
4.1 Comparisons with State-of-the-Art Methods
We compare ObjectClear with open-source state-of-the-art methods in both image inpainting and object removal. The inpainting methods include SDXL-INP [1], PowerPaint [43], and BrushNet [12]. The object removal methods include CLIP-Away [7], DesignEdit [10], RORem [15], FreeCompose [5], and Attentive Eraser [31]. Following MULAN [32], we use the text prompt “an empty scene” for the text-guided baselines when necessary.
Quantitative Evaluation. Unlike our ObjectClear, which explicitly handles both the target object and its associated effects, most baseline methods operate only within the masked regions, overlooking surrounding areas that are visually affected by the object but not explicitly included in the mask. To ensure a fair and comprehensive comparison, we evaluate all methods under two mask conditions: (1) the input mask covers only the target object, and (2) the input mask covers both the target object and its visual effects. These are referred as Object Mask and Object-Effect Mask settings in Table 1, respectively. As shown in Table 1, ObjectClear achieves state-of-the-art performance across all test sets and metrics. Notably, even when using only the object mask, ObjectClear surpasses methods that rely on both object and effect masks. In particular, it achieves a significant advantage in the PSNR-BG metric, highlighting its superior ability to preserve background consistency with the input.
Qualitative Evaluation. Qualitative results are shown in Fig. 4 and Fig. 5. Generation-based inpainting approaches, such as SDXL-INP [24], BrushNet [12], and PowerPaint [43], often generate new objects within the masked regions while fail to remove the effects of the removed objects. In contrast, previous object removal methods, such as Attentive Eraser [31], DesignEdit [10], and RORem [15] demonstrate strong performance in removing the target object itself, but stil fail to eliminate the associated effects. Thanks to our carefully curated training data and the proposed object-effect attention mechanism, ObjectClear effectively removes not only the target object but also its associated shadows and reflections. Moreover, as shown in the last two rows of Fig. 5, ObjectClear effectively handles complex cases involving multiple mutually occluding objects.
Table 2: Quantitative Results of Ablation Studies. Based on our camera-captured data (CC Data), our object removal performance improves progressively by adding the object-effect mask loss ($\mathcal{L}_{mask}$) and simulated training data (Sim. Data). Further improvement is achieved through the proposed Attention-Guided Fusion (AG Fusion).
[…]
(a) 27.29 27.96 0.0910 0.0247
(b) 27.56 28.37 0.0845 0.0217
(c) 28.00 28.81 0.0818 0.0190
(d) 32.98 35.14 0.0352 0.0099
4.2 Ablation Studies
The Effectiveness of the Object-Effect Mask Loss $\mathcal{L}_{mask}$. The object-effect mask loss is designed to explicitly guide ObjectClear to attend to both the target object and its associated effect regions, leading to more thorough and natural removal results. As shown in the chart on the left of Fig. 6, the predicted object-effect attention map achieves a high recall of 0.97. Besides, the attention map in the last column of Fig. 6 on the right, which is from the final denoising step (Step 20), effectively reveals not only the object’s shadow (marked with a red box), but also the object’s reflection in the mirror (marked with a green box), altogether leading to complete and visually coherent object removal results. From Table 2 (a) to (b), improvements can be seen across all metrics with $\mathcal{L}_{mask}$ added, indicating its effectiveness in improving object removal accuracy.
The Effectiveness of Simulated Data. The simulated data effectively enhances the background diversity and scenario complexity of the training set, leading to notable improvements in object removal performance and better background preserving, as shown in Table 2(c). Our carefully designed simulation pipeline, especially the multi-object data synthesis, enables ObjectClear to handle challenging cases involving multiple mutually occluding objects, as demonstrated in the last two rows of Fig. 5. More visualized results can be found in the appendix.
Effectiveness of the Attention-Guided Fusion. The attention-guided fusion operation is designed to preserve background consistency during object removal, a common challenge in generation-based methods such as ObjectDrop [34] and RORem [15], as illustrated in Fig. 7. As the denoising process progresses, the proposed object-effect attention yields increasingly accurate attention maps (Fig.6), which we leverage to guide the fusion of background images. This strategy significantly improves background preservation, as evidenced by notable gains in PSNR-BG. Consequently, we observe a substantial boost in overall performance, as reflected in the quantitative results from (c) to (d) in Table 2, where PSNR improves markedly from $28.00$ to $32.98$.
Figure 7: Effectiveness of Attention-Guided Fusion. The difference maps visualize the pixel-wise differences between the input and output images. RORem [15] fails to remove the reflection of the balloon and introduces changes to the background, resulting in relatively high difference values in the background area and low values in the reflection region. In contrast, both ObjectDrop [34] and our ObjectClear successfully remove the reflection. Notably, ObjectClear better preserves the background, thanks to our predicted object-effect map and proposed attention-guided fusion strategy.
4.3 More Applications
ObjectClear can be flexibly extended to various applications, such as object insertion and movement. We also utilize our OBER dataset to train an insertion network that is able to generate natural object effects. Please refer to the appendix for more details. As shown in Fig. 8, even when only the target objects are specified for insertion and movement, ObjectClear is capable of generating plausible and natural shadows and reflections accordingly.
Figure 8: Object Insertion and Movement. In addition to accurately inserting or repositioning objects, our ObjectClear also generates plausible and natural shadows and reflections accordingly.
5 Conclusion
We introduce ObjectClear, a practical framework for object removal that ensures high-quality object-effect removal and background-preserving results across diverse real-world scenarios. Our method employs an Object-Effect Attention (OEA) mechanism to adaptively focus on removal regions, including both the object and its effects, coupled with Attention-Guided Fusion (AGF) to preserve background details accurately. We also propose OBER, a large-scale and diverse hybrid dataset combining camera-captured and simulated data. With the carefully designed mechanism and curated dataset, ObjectClear achieves robust and superior performance, effectively addressing existing challenges in object removal tasks. In addition, ObjectClear can be easily adapted to various applications, such as object insertion and object movement, and consistently achieves superior performance. These advancements position ObjectClear as a reliable solution for practical image editing, laying a solid foundation for future research in controllable image processing through local editable regions.
Appendix
In this supplementary material, we provide additional discussions and results to complete the main paper. In Sec. A, we present further details on data augmentation during training and other extended applications. In Sec. B, we provide more details of the proposed OBER dataset, including statistics and some examples for demonstration. In Sec. C, we report additional results, such as more results on ablation studies, a user study, experiments with user strokes, more results of object insertion and movement, and additional comparisons on in-the-wild data.
A More Details of ObjectClear
A.1 Training Augmentations
During training, we apply on-the-fly random cropping to enable the network to learn object removal across varying object sizes. We also apply color augmentation and random flipping to enhance model robustness. To improve the model’s robustness to the estimated or user-provided coarse mask, several previous methods [34, 11, 33] have introduced mask augmentation techniques. In line with these approaches, we also apply dilation and erosion to the input mask during training. Different from previous practice, our method employs an object-aware dilation and erosion strategy, where the dilation and erosion kernel size is adaptively determined based on the size of the object. As demonstrated in the qualitative results in Sec. C.3, our method effectively handles coarse user-drawn masks by implicitly completing and refining them, showcasing strong robustness to imprecise inputs.
A.2 More Details for Other Applications
In this section, we further discuss implementation details of other applications, including object insertion and object movement.
Object Insertion. The insertion network also leverages the OBER dataset for training and adopts the same architecture as ObjectClear (Fig. 3 of the main paper), thus receiving the input tuple of $<z_{t},I_{in},M_{o},c>$ and supervised by $I_{GT}$ (for output image) and $M_{fg}$ (for object-effect attention). However, $I_{in}$, $I_{GT}$, and $c$ are constructed in an “reverse” way compared with the removal network.
While the ground truth $I_{GT}$ is the original image with the object accompanied by natural effects, $I_{in}$ is the one with the object simply copied and pasted on the background. Specifically, to generate $I_{in}$, we first extract the object from $I_{GT}$ with the object mask $M_{o}$ only, and then we paste it onto the corresponding background image. In addition to the image pair, the input text $c$ is also reversed to “insert the instance of”. Notably, the object-effect attention map is still supervised by $M_{fg}$. In this context, $M_{fg}$ refers to the mask that covers both the object and its generated effects (e.g., shadows or reflections).
During inference, we directly paste an object (without effects) onto the background scene, along with its object mask $M_{o}$, which forms the input image $I_{in}$. The network then generates the output image where both the object and its generated effects are harmoniously inserted. Since the insertion network also integrates the object-effect attention, we can also apply Attention-Guided Fusion to preserve the background fidelity while generating realistic object effects.
Object Movement. To enable object movement, we combine the ObjectClear and insertion network introduced above. Specifically, we first apply ObjectClear to remove the target object along with its associated effects (e.g., shadows or reflections), resulting in a clean object-free background. The object is then extracted using its given object mask. Users are allowed to specify a new location and optionally adjust the object scale before the object is re-inserted into the clean background. With our insertion network, the object is harmonized with the new context by generating realistic effects. This two-stage approach supports controllable object movement while ensuring visual realism and consistency.
B More Details of OBER Dataset
Table 3: Dataset Overview. The OBER dataset consists of camera-captured and simulated training data, as well as two testing sets: OBER-Test (w/ ground truth) and OBER-Wild (w/o ground truth). All data subsets are annotated with object masks and object-effect masks.
Properties Training Set Testing Set
Camera-Captured Simulated OBER-Test OBER-Wild
#Image Pairs 2,715 10,000 160 300
w/ Ground Truth ×
w/ Object Mask
w/ Object-Effect Mask
Table 4: Comparison of OBER Dataset with Existing Datasets. * indicates that the Counterfactual dataset is not publicly available.
Description RORD [29] MULAN [32] DESOBAv2 [18] Counterfactual [34] OBER (ours)
Tasks Object Removal × ×
Effect Removal × ×
Object-Effect Removal × ×
Annotations Object Mask ×
Effect Mask × × ×
Object-Effect Mask ✓(Coarse) × ×
RGBA Objects × × × ×
Multi Objects ×
Effect Types Shadow × Shadow Shadow Shadow & Reflection
Camera-Captured GT × ×
Overview. Table 3 provides an overview of our OBER dataset. The training set consists of 2,715 camera-captured image pairs and 10,000 simulated pairs, all annotated with object masks and object-effect masks. For evaluation, we provide two test subsets: OBER-Test (includes 160 pairs with ground truth) and OBER-Wild (includes 300 pairs without ground truth), where object masks and object-effect masks are also available. In addition, we showcase some samples from our OBER dataset in Fig. 9, including camera-captured data and simulated data, alongside with their annotations such as object masks, object-effect masks, and RGBA object foregrounds.
Reflection Pair Simulation. Thanks to the strong priors of generative models, we observed that the model trained without reflection data was still able to remove some simple reflection effects, a finding consistent with the conclusions drawn in ObjectDrop [34]. However, the model often struggled with more challenging reflection removal cases, such as when the object and its reflection were spatially separated or when the reflection was heavily distorted by surface ripples. To address this limitation, we adopted a human-in-the-loop strategy to collect paired reflection data. Specifically, we used our trained model to perform inference on 200 real-world reflection images and manually selected 50 high-quality results (as shown in Fig. 9(b)), which were then added as an important supplement to the training data. We found that even a small amount of high-quality reflection data pairs could considerably improve the model’s ability to generalize across diverse reflection scenarios.
Comparison with Existing Datasets. We compare our OBER dataset with existing datasets, including those focused on object removal (RORD [29], MULAN [32], Counterfactual Dataset [34]) and one on shadow removal (DESOBA-v2 [18]). As summarized in Table 4, our OBER is the only dataset that includes all three types of mask annotations and RGBA foreground objects, enabling a wide range of tasks such as object removal, effect removal, and joint object-effect removal.
While RORD [29] is among the first to provide object-effect masks, those masks are coarse without separate annotations for objects and their effects. As a result, it cannot support independent removal of either the object or the effect, but only their joint removal. In contrast, our OBER dataset provides precise and separate masks for objects and their effects, enabling more flexible and fine-grained removal tasks. In addition, MULAN [32] focuses solely on object removal without addressing associated effects such as shadows or reflections. DESOBA-v2 [18] targets only effect removal and thus does not handle the objects. Moreover, the ground truth images in both datasets are not captured by cameras but are instead synthesized using generative models, which may limit their realism. The Counterfactual Dataset proposed in ObjectDrop [34] is not publicly available, and it only includes object masks without object-effect masks. It also lacks foreground RGBA data for objects and associated effects, limiting its extensibility and applicability to more advanced object editing tasks.
Figure 9: Samples from our OBER Dataset. We provide some samples from our OBER dataset including both camera-captured data and simulated data. While all data from both categories are annotated with fine-grained object masks and object-effect masks, we also extract RGBA foreground objects with associated effects from camera-captured data (as discussed in Sec. 3.1 in the main paper), which could be used to construct realistic simulated samples. In the simulated samples, we not only include the shadow effects but also the reflection effects, which are rarely included in other datasets (Table 4), thanks to our reflection pair simulation strategy detailed in Sec. B.
C More Results
C.1 Results for Ablation Study
Effectiveness of $\mathcal{L}_{mask}$. Thanks to the rich annotations in our OBER dataset, the object-effect mask enables a supervision loss, denoted as $\mathcal{L}_{mask}$, which guides the cross-attention layers to focus on both the object and its associated effects, while preserving background textures. This supervision facilitates a decoupled optimization of object removal and background reconstruction. As shown in Fig. 10, with the mask loss $\mathcal{L}_{mask}$, the network can adaptively identify the object-effect regions to be removed, as reflected in the attention mask (shown in yellow boxes). This leads to more accurate and complete removal of objects and their shadow effects, without mistakenly erasing unrelated background content.
Figure 10: Effectiveness of $\mathcal{L}_{mask}$. It could be observed that when training without the $\mathcal{L}_{mask}$ supervision, the model struggles to accurately remove the target object and its effects, leading to mistakenly erasing unrelated background object (right) or effects (left). In contrast, when supervised with $\mathcal{L}_{mask}$, the attention mask (shown in yellow boxes) could identify the removal regions well, thus leading to more accurate and complete removal results.
Effectiveness of Simulated Data. To balance data realism and scalability, in addition to the camera-captured data, we scale up our OBER dataset with a carefully designed simulation pipeline. Our simulation data is generated by compositing the foreground RGBA object (extracted from camera-captured data) onto diverse backgrounds. In particular, the simulation of multi-object compositions leads to notable improvements in object removal robustness when mutually occluding objects exist, as shown in Fig. 11 (left). Furthermore, our simulated reflection image pairs (discussed in Sec B) greatly improve the model capability of removing reflections, including challenging cases with significant distortion due to water surface ripples, as shown in Fig. 11 (right).
Figure 11: Effectiveness of Simulated Data. Since our simulation data includes multi-object compositions, training with such data enables the model to accurately remove the target object and its associated effects while preserving unrelated object effects (left). In addition, adding the reflection data pairs during training greatly enhances the model capability of removing reflections, including challenging cases with significant distortion due to water surface ripples (right).
Effectiveness of Attention-Guided Fusion. The attention map supervised by $\mathcal{L}_{mask}$ supports the Attention-Guided Fusion (AG Fusion) strategy during inference. It helps to blend the generated image with the original input via a copy-and-paste operation, where pixels within the object-effect region are taken from the generated result, and the rest are preserved from the original image. Such practice effectively reduces undesired background detail changes caused by VAE reconstruction errors and the diffusion process, thereby greatly preserving the background fidelity. In Fig. 12, we visualize the background detail changes by showing the difference maps between the generated image and the corresponding input, where a clear improvement on background preservation could be observed.
Figure 12: Effectiveness of Attention-Guided Fusion (AG Fusion). We visualize the background detail changes by showing the difference maps between the generated image and the corresponding input. In ideal cases, the difference maps should show large values only in the object removal regions, including the object and its effects, while showing little difference in the unrelated background. It could be observed that when AG Fusion is not employed, although the difference in object removal regions is significant, the difference in the unrelated background is also noticeable. However, with the AG Fusion strategy, we almost eliminate the undesirable background difference thanks to the accurate attention map predicted.
C.2 User Study
To enable a more comprehensive evaluation, we conducted a user study on object removal results for in-the-wild images. We compared ObjectClear with two representative state-of-the-art methods: PowerPaint [43] and RORem [15]. For a fair comparison, we evaluated each baseline method under two settings: (1) conditioned on the object mask and (2) conditioned on the object-effect mask. In contrast, all ObjectClear results were generated using only the object mask, and we compared our outputs against both settings of the baselines.
We invite a total of 30 participants for this user study. Each volunteer was presented 40 randomly selected image quadruples, consisting of: an input image, two results from a baseline method under different mask conditions, and our result. Participants were asked to select the best removal result based on two criteria: the realism of the object region and the preservation of background details.
As summarized in Table 5, ObjectClear outperforms the baselines under both mask settings. Notably, although baseline methods benefited from access to object-effect masks, our ObjectClear still won more user preference with the object mask only.
Table 5: User Study. Our ObjectClear is preferred by human voters over other two representative state-of-the-art methods (PowerPaint [43] and RORem [15]) under both settings. Although the compared methods won more user preference given the object-effect mask instead of the object mask, our ObjectClear still received far more user preference with the object mask only.
Methods PowerPaint (Object Mask) [43] PowerPaint (Object+Effect Mask) [43] ObjectClear (Ours)
User Preference 2.68% 7.03% 90.29%
Methods RORem (Object Mask) [15] RORem (Object+Effect Mask) [15] ObjectClear (Ours)
User Preference 3.51% 11.80% 84.69%
C.3 Results with User Strokes
In practical applications, users often interact with visual systems through imprecise or casually drawn inputs, such as rough scribbles or incomplete masks. These inputs may vary significantly in shape, location, and accuracy. Therefore, it is essential for a robust object removal network to effectively process such arbitrary mask inputs without relying on carefully crafted annotations. Benefiting from our mask augmentation strategy and object-effect attention mechanism, our network demonstrates strong robustness to diverse mask inputs. In this subsection, we simulate user strokes and feed them into the network along with the images. The resulting outputs and attention maps show that our network can accurately identify and attend to the object and its associated effects, even with imprecise masks, as illustrated in Fig. 13.
Figure 13: Results with User Strokes. We simulate user strokes and feed them into the network along with the images. The resulting outputs and attention maps show that our network can accurately identify and attend to the object and its associated effects, even with imprecise masks.
C.4 More Comparisons on In-the-wild Data
In this subsection, we compare ObjectClear with state-of-the-art methods, including SDXL-INP [1], PowerPaint [43], Attentive Eraser [31], DesignEdit [10], and RORem [15], on in-the-wild data. To ensure a fair comparison, we consider two settings: (1) All methods use the same object mask; (2) While other methods use finely annotated object-effect masks, our ObjectClear still uses the object mask. The results under these two settings are shown in Fig. 14 and Fig. 15 separately. Under the first setting, when given the object mask only, our ObjectClear effectively removes shadow and reflection associated with the target object, while all the other methods fail to remove those effects and tend to hallucinate undesirable content or create artifacts. Under the second setting, while other methods do remove the shadow thanks to the given effect region, they also alter or remove the original background content in an undesirable way. In contrast, our ObjectClear removes the object effect accurately with the object mask only while preserving the background with high fidelity.
Figure 14: Object Removal on OBER-Wild - Condition on Object Mask. Given the object mask only, our ObjectClear effectively removes shadow and reflection associated with the target object, while all the other methods fail to remove those effects and tend to hallucinate undesirable content or create artifacts.
Figure 15: Object Removal on OBER-Wild - Comparison between methods conditioned on Object-Effect Mask (others) and Object Mask (ours). Since existing methods struggle to remove object effects when provided with only the object mask, we supply them with our annotated object-effect masks for a fair comparison. Although these methods are able to remove shadows with the additional effect region, they often introduce undesirable changes to the original background. In contrast, ObjectClear effectively removes the object and its associated effects using only the object mask, while preserving the background content with high fidelity.
C.5 Limitations
While ObjectClear exhibits strong performance in removing objects and their associated effects, it still faces challenges in highly complex scenarios. Specifically, (1) when multiple shadows heavily overlap, the model struggles to accurately disentangle and localize the effect corresponding to the target object; and (2) in cases involving extremely large and long shadow regions, ObjectClear may fail to completely remove the effect, occasionally leaving behind residual artifacts. Enhancing the disentanglement of complex effect regions and improving the removal of large-scale effects remain important directions for future research.

### v2

@@
-ObjectClear: Complete Object Removal via Object-Effect Attention
+Precise Object and Effect Removal with Adaptive Target-Aware Attention
 Abstract
-Object removal requires eliminating not only the target object but also its effects, such as shadows and reflections.
-However, diffusion-based inpainting methods often produce artifacts, hallucinate content, alter background, and struggle to remove object effects accurately.
-To address this challenge, we introduce a new dataset for OBject-Effect Removal, named OBER, which provides paired images with and without object effects, along with precise masks for both objects and their associated visual artifacts.
-The dataset comprises high-quality captured and simulated data, covering diverse object categories and complex multi-object scenes.
-Building on OBER, we propose a novel framework, ObjectClear, which incorporates an object-effect attention mechanism to guide the model toward the foreground removal regions by learning attention masks, effectively decoupling foreground removal from background reconstruction.
-Furthermore, the predicted attention map enables an attention-guided fusion strategy during inference, greatly preserving background details.
-Extensive experiments demonstrate that ObjectClear outperforms existing methods, achieving improved object-effect removal quality and background fidelity, especially in complex scenarios.
-footnotetext: Corresponding author.
+Object removal requires eliminating not only the target object but also its associated visual effects such as shadows and reflections.
+However, diffusion-based inpainting and removal methods often introduce artifacts, hallucinate contents, alter background, and struggle to remove object effects accurately.
+To address these challenges, we propose ObjectClear, a novel framework that decouples foreground removal from background reconstruction via an adaptive target-aware attention mechanism.
+This design empowers the model to precisely localize and remove both objects and their effects while maintaining high background fidelity.
+Moreover, the learned attention maps are leveraged for an attention-guided fusion strategy during inference, further enhancing visual consistency.
+To facilitate the training and evaluation, we construct OBER, a large-scale dataset for OBject-Effect Removal, which provides paired images with and without object-effects, along with precise masks for both objects and their effects.
+The dataset comprises high-quality captured and simulated data, covering diverse objects, effects, and complex multi-object scenes.
+Extensive experiments demonstrate that ObjectClear outperforms prior methods, achieving superior object-effect removal quality and background fidelity, especially in challenging scenarios.
+Figure 1: Object Removal Comparison.
+Given an object mask, prior methods often leave residual artifacts or hallucinate undesirable content, change background, and usually fail to remove associated effects such as shadows and reflections.
+In contrast, our ObjectClear precisely eliminates both the object and its associated effects, achieving seamless object removal results even in challenging cases.
+footnotetext: Equal contribution  Corresponding author
 1 Introduction
-Figure 1: Object Removal Comparison.
-Given an object mask, prior methods often leave residual artifacts or hallucinate undesirable content, change background, and typically fail to remove associated effects such as shadows and reflections.
-In contrast, our ObjectClear precisely eliminates both the object and its associated effects, achieving seamless object removal results even in challenging cases.
-Recent advances in generative models [27, 24, 21], such as GPT-4o, have shown strong capabilities in image editing and are widely adopted in real-world applications.
-Among these tasks, object removal has become a key task, allowing users to erase unwanted content from images.
-However, seamlessly removing both the object and its effects (e.g., shadows and reflections) while preserving the background remains a challenging problem.
-While diffusion-based methods [1, 43, 12, 15, 31, 7, 10, 5] have advanced object removal, they still struggle with high-fidelity results.
-As shown in Fig.
-1, existing methods often leave residual artifacts, hallucinate unwanted content, alter backgrounds details, and fail to remove object effects.
-This is partly because the input mask fully occludes the target object, preventing models from perceiving its surrounding context and associated effects.
-In addition, the lack of large-scale and public datasets for object-effect removal also limits further progress in the field.
-Existing object removal datasets can be categorized into simulated and camera-captured data.
+Recent advances in generative models [30, 26, 23], have demonstrated strong capabilities in image editing and are widely adopted in real-world applications.
+Among these tasks, object removal has become a key application, enabling users to erase unwanted content from images.
+However, seamlessly removing both the object and its associated visual effects (e.g., shadows and reflections) while preserving the background remains a challenging problem, as it requires precisely localizing the object together with its effects, restoring the removed regions with background visual consistency, and maintaining the rest of the image unchanged.
+With the rapid development of diffusion-based image generation models [8, 26, 14, 15] and object segmentation techniques [13, 28], combining diffusion-based image generators with target object masks has become the mainstream paradigm for object removal [26, 48, 12, 18, 34, 6, 10, 4].
+However, as shown in Figs.
+1 and 5, existing methods often fail to remove the visual effects of target objects, generate undesired objects in the removed areas, or unintentionally modify the background (see Fig.
+7).
+These limitations mainly stem from the lack of explicit modeling of the correlation between the target object and its associated effects, as well as the absence of effective constraints to guide the generative model’s attention toward the removal regions.
+To address these challenges, we propose ObjectClear, a novel framework designed to achieve precise and complete removal of both objects and their visual effects while preserving background fidelity.
+To enhance the modeling of object–effect regions, we introduce an Adaptive Target-Aware Attention (ATA) mechanism that learns to localize objects and their associated effects under the supervision of object–effect masks.
+By integrating text–image embeddings, ATA effectively decouples foreground removal from background reconstruction, enabling the model to remove targets accurately and suppress undesired object generation within the removed regions.
+During inference, we further introduce an Attention-Guided Fusion (AGF) module that leverages ATA-predicted object–effect attention maps to perform adaptive input–output fusion, preserving fine-grained background details including color and texture.
+In addition, we introduce a Spatially-Varying Denoising Strength (SVDS) strategy to address incomplete removal and color inconsistency caused by uniform denoising strength.
+Together, these designs empower ObjectClear to deliver precise and high-fidelity effect-object removal.
+To train ObjectClear and fully realize its potential, a dedicated dataset containing paired images with and without objects, along with precise masks for both objects and their effects, is essential.
+Unfortunately, to the best of our knowledge, no such large-scale training dataset is publicly available.
+Current object removal datasets can be categorized into simulated and camera-captured data.
 (1) Simulated data.
@@
 (2) Camera-captured data.
-Some works [29, 33] leverage fixed-viewpoint videos to extract paired frames, but this constrains the foreground types to moving objects.
-Others [34, 39, 36] capture image pairs before and after object removal, but these datasets are often not publicly available and small in scale due to the high costs.
-To bridge this gap, we introduce OBER, a hybrid dataset for OBject-Effect Removal, which combines both camera-captured and simulated data with diverse foreground objects, background scenes, and object effects.
+Some works [31, 37] extract paired frames from fixed-viewpoint videos, but this restricts the foreground to moving objects only and makes it difficult to maintain background consistency between paired samples.
+Others [38, 42, 39] capture image pairs before and after object removal, but these datasets remain small in scale due to their high collection cost and are not publicly released.
+To overcome these limitations, we present OBER, a hybrid dataset for OBject-Effect Removal, which combines both camera-captured and simulated data with diverse foreground objects, background scenes, and object effects (e.g., indoor/outdoor shadows and reflections).
 For the camera-captured data, we annotate precise object and object-effect masks, which serve as critical supervision during training.
@@
 By combining the realism of captured data with the scalability of simulation, OBER offers a high-quality, large-scale, and diverse dataset with a total of 12,715 training samples.
-In addition, we introduce two new benchmarks: OBER-Test (w/ ground truth) and OBER-Wild (w/o ground truth), to support future research.
-Built upon OBER, we propose a novel framework, ObjectClear, which introduces an Object-Effect Attention (OEA) mechanism that adaptively focuses on the foreground removal region, including both the object and its associated effects, by supervising cross-attention maps with object-effect masks.
-The OEA module effectively decouples foreground removal from background reconstruction, enabling more precise and complete object elimination.
-In addition, the predicted attention map supports an Attention-Guided Fusion (AGF) strategy during inference, helping to preserve background details, as shown in Fig.
-7.
-This design enables the model to localize editable regions adaptively and offers a potential solution to the long-standing fidelity issues in diffusion-based image editing.
-Our contributions are summarized as follows:
-  • We propose OBER, a high-quality and large-scale hybrid dataset for object removal, featuring diverse objects, fine-grained annotations of object-effect masks, and complex multi-object scenarios across both simulated and camera-captured settings.
-  • We introduce a novel framework, ObjectClear, which incorporates an Object-Effect Attention mechanism that adaptively focuses on foreground removal regions and enables an Attention-Guided Fusion strategy, improving object removal quality and background fidelity.
-  • We present two new benchmarks, OBER-Test and OBER-Wild, providing a foundation for future research.
-Our method achieves superior performance on all benchmarks, outperforming existing approaches in terms of both quantitative metrics and visual quality.
+In addition, we introduce two new benchmarks: OBER-Test and OBER-Wild, to support future research.
+Our contributions are summarized as follows: (1) We introduce a novel framework, ObjectClear, which incorporates an Adaptive Target-Aware Attention (ATA) mechanism that adaptively focuses on foreground removal regions, together with an Attention-Guided Fusion (AGF) and a Spatially-Varying Denoising Strength (SVDS) strategy, thereby improving object removal quality and background fidelity.
+(2) We propose OBER, a high-quality and large-scale hybrid dataset for object-effect removal, featuring diverse objects, fine-grained annotations of object-effect masks, and complex scenarios across both simulated and camera-captured settings.
+(3) ObjectClear achieves superior performance on both synthetic and real-world benchmarks, demonstrating clear advantages in quantitative results and visual quality.
 2 Related Work
+Figure 2: Dataset Construction Pipeline of OBER.
+The dataset combines both camera-captured and simulated data, featuring diverse foreground objects and background scenes.
+It provides rich annotations, including object masks, object-effect masks, transparent RGBA object layers, and complex multi-object scenarios for training and evaluation.
 Image Inpainting.
@@
 Early approaches predominantly adopt generative adversarial networks (GANs) [7], but often suffer from limited realism and diversity [19, 25, 29, 43].
-With the rapid advances in diffusion models [28, 30, 9, 24], many methods [2, 20, 21, 43, 12, 38] have begun leveraging their strong generative priors to synthesize high-fidelity content, achieving state-of-the-art results in image inpainting.
+With the rapid advances in diffusion models [30, 33, 9, 26], many methods [1, 22, 23, 48, 12, 41, 36] have begun leveraging their strong generative priors to synthesize high-fidelity content, achieving state-of-the-art results in image inpainting.
 In this work, we adapt the SDXL-Inpainting model [26] for photorealistic completion.
-However, despite their strong generative capabilities, existing inpainting models often lack awareness of object-induced effects (e.g., shadows and reflections), leading to incomplete or visually inconsistent object removal results.
+However, despite their strong generative capabilities, existing inpainting models often lack awareness of object effects (e.g., shadows and reflections), leading to incomplete or inconsistent object removal results.
 Object Removal.
-Object removal is a specialized form of image inpainting that requires accounting for object effects to ensure complete and realistic removal, which diffusion-based models [34, 11, 19, 39, 15, 4, 43, 7, 5, 36, 10, 31, 33] currently dominate.
-A common strategy involves curating high-quality triplet datasets.
-For instance, ObjectDrop [34] builds a real-world dataset by having photographers capture the same scene before and after the removal of a single object.
-However, its limited scale and lack of public availability hinder broader adoption.
-To address data scalability, methods like SmartEraser [11] and Erase Diffusion [19] rely on synthetic datasets using segmentation or matting techniques to extract foreground objects, but these typically lack annotations of object effects, limiting the models’ ability to remove shadows or reflections.
-To enhance realism, methods such as LayerDecomp [36] and OmniPaint [39] create costly camera-captured data.
-OmniPaint [39] takes one step further to annotate unlabeled images by a model trained on small-scale real data, while RORem [15] involves human annotators to ensure the quality of annotated data.
-In parallel, works like RORD [29] and OmniEraser [33] scale data generation by mining realistic video frames with fixed viewpoints, selectively pairing frames with and without target objects while preserving natural object effects.
+Object removal is a specialized branch of image inpainting that often requires explicit consideration of object effects (e.g., shadows and reflections) to achieve complete removal, a problem increasingly dominated by diffusion-based methods [38, 11, 21, 42, 18, 3, 48, 6, 4, 39, 10, 34, 37, 47], and recently extended to the video domain [46, 16, 24, 32].
+A common strategy is to curate triplet datasets.
+ObjectDrop [38] builds a real-world dataset by capturing the same scene before and after object removal, but its limited scale and non-public availability hinder broader adoption.
+To improve data scalability, methods like SmartEraser [11] and EraseDiffusion [21] rely on synthetic datasets generated by segmentation or matting methods to extract foreground objects, but these datasets typically lack annotations of object effects, limiting the ability to remove shadows or reflections.
+OmniPaint [42] further expands its dataset by automatically annotating unlabeled images using a model trained on small-scale real data, while RORem [18] also incorporates human annotators to ensure annotation quality.
+In parallel, works like RORD [31] and OmniEraser [37] scale data generation by mining realistic video frames with fixed viewpoints, selectively pairing frames w/ and w/o target objects while preserving natural object effects.
+GeoRemover [47] leverages depth geometry to achieve precise object removal.
 To eliminate reliance on curated datasets altogether, other methods opt for test-time optimization [4, 34, 10] or providing plug-and-play solutions [6].
-However, all these models rely on implicit learning of object effects without explicitly modeling effect maps, making it difficult to ensure that no unnecessary background changes occur.
+However, these models rely on implicitly learning object effects without explicitly modeling effect maps, which makes it difficult to maintain background consistency.
+Figure 3: The Framework of ObjectClear.
+Given an input image and a target object mask, ObjectClear employs an Adaptive Target-Aware Attention mechanism to guide the model toward foreground removal regions by learning attention masks.
+The predicted mask further enables an Attention-Guided Fusion strategy during inference, which substantially preserves background details.
 3 Methodology
-To achieve accurate removal of target objects along with their associated visual effects, such as shadow and reflection, we propose a comprehensive framework consisting of a data curation pipeline (OBER Dataset) and a dedicated object removal model (ObjectClear).
-These two components are detailed in Sec.
-3.1 and Sec.
-3.2, respectively.
+To achieve accurate removal of target objects and their associated effects, we propose a comprehensive framework with a data curation pipeline (OBER Dataset, Sec.
+3.1) and a dedicated object removal model (ObjectClear, Sec.
+3.2).
 3.1 OBER Dataset
-ObjectDrop [34] shows that even small-scale datasets, when captured with careful camera control over scene consistency, can significantly enhance model generalizability in object removal tasks.
-However, due to the lack of public access to such datasets and the high cost of real-world data collection, subsequent works often face limitations in dataset scale.
-To address this challenge, we introduce the OBER dataset (OBject-Effect Removal), designed to balance data realism and scalability.
-As illustrated in Fig.
-2, OBER is a hybrid dataset consisting of two parts: (1) a small set of camera-captured images adhering to physical realism, and (2) a larger set of simulated images generated by compositing foreground objects, extracted from the real data in (1), onto diverse background scenes.
-Figure 2: Dataset Construction Pipeline of OBER.
-The dataset combines both camera-captured and simulated data, featuring diverse foreground objects and background scenes.
-It provides rich annotations, including object masks, object-effect masks, transparent RGBA object layers, and complex multi-object scenarios for training and evaluation.
+To learn an object–effect removal model, a high-quality dataset with paired images and precise masks for both objects and their effects is required.
+However, to our knowledge, no such large-scale dataset is publicly available.
+To this end, we introduce the OBER dataset (OBject-Effect Removal), designed to balance data realism and scalability.
+As shown in Fig.
+2, OBER is a hybrid dataset consisting of two parts: (1) a small set (2,878) of camera-captured images adhering to physical realism, and (2) a larger set (10,000) of simulated images generated by compositing foreground objects, extracted from (1), onto diverse background scenes.
 Camera-Captured Data.
@@
 For each pair, the image with the object is used as the input $I_{in}$, and the image without the object serves as the ground truth $I_{gt}$.
-In total, we collected 2,875 such counterfactual pairs with 2,715 for training and 160 for testing.
+In total, we collected 2,878 such counterfactual pairs with 2,715 for training and 163 for testing.
 Our dataset encompasses a wide variety of everyday objects commonly found in indoor (e.g., benches, cups, dolls) and outdoor scenes (e.g., pedestrians, vehicles).
@@
 Pixels with differences above a predefined threshold are regarded as part of the object-effect mask.
-This object-effect mask provides crucial supervision during training, allowing the network to adaptively learn to focus on the target removal regions, including both the object itself and its effects.
+This object-effect mask provides crucial supervision during training, allowing the network to adaptively learn to focus on the target removal regions, including both the object itself and its associated effects.
 Simulated Data.
@@
 2(b)).
-$\alpha(p)=\begin{cases}0,&\text{if pixel }p\in\text{background area $\overline{M_{o}}\cap\overline{M_{e}}$}\\ 1,&\text{if pixel }p\in\text{object area $M_{o}$}\\ (I_{\text{gt}}-I_{\text{in}})/(I_{\text{gt}}+\varepsilon),&\text{if pixel }p\in\text{effect area $M_{e}$}\\ \end{cases}$ (1)
+$\alpha(p)=\begin{cases}0,&\text{if }p\in\text{$\overline{M_{o}}\cap\overline{M_{e}}$}\\ 1,&\text{if }p\in\text{$M_{o}$}\\ (I_{\text{gt}}-I_{\text{in}})/(I_{\text{gt}}+\varepsilon),&\text{if }p\in\text{$M_{e}$}\\ \end{cases},$ (1)
+where $M_{o}$ and $M_{e}$ are object and effect regions, respectively.
 (3) Blend Objects with Backgrounds.
@@
 2(b).
+These designs ensure physically plausible placement and consistent lighting, reducing the domain gap between simulated and real data.
+Synthetic data also overcomes challenges in real-world collection, such as limited scale and complex cases (e.g., multi-object occlusions, reflections), while its controllability expands coverage and diversity.
 In total, we generate 10,000 composite images, significantly enriching the diversity and scalability of our dataset.
@@
 3.2 ObjectClear
-Figure 3: The Framework of ObjectClear.
-Given an input image and a target object mask, ObjectClear employs an Object-Effect Attention mechanism to guide the model toward foreground removal regions by learning attention masks.
-The predicted mask further enables an Attention-Guided Fusion strategy during inference, which substantially preserves background details.
 Our ObjectClear is built upon SDXL-Inpainting [26].
-While SDXL-Inpainting uses a noise map $z_{t}$, a masked image $I_{m}$, a corresponding mask $M_{o}$, and a text prompt $c$ as inputs, we feed the original image $I_{in}$ instead of $I_{m}$ into the model, and our inputs are expressed as a tuple $<z_{t},I_{in},M_{o},c>$.
+While SDXL-Inpainting uses a noise map $z_{t}$, a masked image $I_{m}$, a corresponding mask $M_{o}$, and a text prompt $c$ as inputs, we feed the original image $I_{in}$ instead of $I_{m}$ into the model, and our inputs are expressed as a tuple $z_{t},I_{in},M_{o},c>$.
 This design encourages the model to better attend to the effect of the target object by leveraging its visual features.
-Moreover, it facilitates more effective utilization of background information behind the object in cases involving moving transparent objects, such as glass cups.
-Apart from the input design, we further introduce an Object-Effect Attention Mechanism and an Attention-Guided Fusion Strategy to explicitly attend object-effect regions and enhance background preservation.
-Detailed descriptions of these components are provided below.
-Object-Effect Attention.
-To enable the model to better attend to both the target object region and its associated visual effect regions, we follow the approach of Li et al.
-[16] by stacking text and visual object information as guidance for the cross-attention layers in the base model.
+Moreover, it facilitates more effective utilization of background information behind the object when transparent objects are to be removed, such as glass cups.
+To achieve precise and complete object removal while enhancing background preservation, we introduce Adaptive Target-Aware Attention (ATA), Attention-Guided Fusion (AGF), and Spatially-Varying Denoising Strength (SVDS) strategy, which explicitly attend to object-effect regions.
+Adaptive Target-Aware Attention.
+To enable the model to better attend to both the target object region and its associated effect regions, we integrate text and object image as an object prompt for the cross-attention layers in the base model.
 Specifically, the text prompt is expressed as “remove the instance of”, and the visual object is obtained by applying a dot product between the input image $I_{in}$ and $M_{o}$, as illustrated in Fig.
@@
 Concretely, we extract the cross-attention maps corresponding to the visual embedding tokens and denote them as $\mathbf{A}$, which we supervise with the annotated foreground object-effect masks $M_{fg}$ from our OBER dataset.
-Following [35, 23], $\mathcal{L}_{\text{mask}}$ is designed to minimize the attention values in the background regions while maximizing those in the foreground.
+$\mathcal{L}_{\text{mask}}$ is designed to minimize the attention values in the background regions while maximizing those in foreground.
 This objective can be formulated as:
 $\mathcal{L}_{mask}=\mathrm{mean}({\mathbf{A}}[1-M_{fg}])-\mathrm{mean}({\mathbf{A}}[M_{fg}])\text{,}$ (2)
-where []delimited-[][$\cdot$][ ⋅ ] denotes indexing operator, and $\mathbf{A}[M_{fg}]$ refers to the attention values within the foreground regions specified by the mask $M_{fg}$.
+where $\cdot$] denotes indexing operator, and $\mathbf{A}[M_{fg}]$ refers to the attention values within the foreground regions specified by the mask $M_{fg}$.
 Attention-Guided Fusion.
-Interestingly, we observe that applying Object-Effect Attention not only improves ObjectClear’s precision in object removal, but also produces attention maps that accurately capture both the object and its effects.
-Based on this observation, we propose Attention-Guided Fusion strategy that utilizes attention maps during inference to seamlessly combine the generated result with the original image.
+Interestingly, we observe that applying Adaptive Target-Aware Attention not only improves ObjectClear’s precision in object removal, but also produces attention maps that accurately capture both the object and its effects.
+Based on this observation, we propose an Attention-Guided Fusion strategy that leverages the predicted attention maps during inference to seamlessly blend the generated result with the original image.
 Specifically, we extract the first-layer cross-attention map corresponding to the object embedding during inference.
-This attention map is then upsampled to match the spatial resolution of the original image, forming a soft estimate of the object-effect region.
-To enhance the spatial smoothness and avoid hard edges, we apply a Gaussian blur to the upsampled attention map, producing a softened object-effect mask.
-This blurred mask is then used to blend the generated image with the original input via a copy-and-paste operation, where pixels within the object-effect region are taken from the generated result, and the rest are preserved from the original image.
-This strategy significantly reduces undesired background changes caused by the diffusion process and VAE reconstruction, while preserving high-fidelity object removal.
-Unlike BrushNet [12], which relies on user-provided object masks to guide blending, our approach adaptively produces the object-effect mask from our Object-Affect Attention.
+This attention map is then upsampled to match the resolution of the original image, forming a soft estimate of the object-effect region.
+To avoid edge artifacts during blending, we apply a Gaussian blur to the upsampled attention map, producing a soft-edged object-effect mask.
+This mask is then used in an alpha-blending operation to fuse the generated image with the original input.
+This strategy significantly reduces undesired background changes introduced by the diffusion denoising process and VAE reconstruction, achieving high-fidelity object removal.
+Unlike BrushNet [12], which relies on user-provided object masks that usually exclude effect regions, our approach leverages precise object-effect masks adaptively generated by Adaptive Target-Aware Attention.
+Spatially-Varying Denoising Strength.
+In diffusion-based image editing, the initial latent for denoising is typically obtained by adding a certain amount of noise to the input image latent.
+The denoising strength (DS) $\mathrm{DS}\in[0,1]$ controls the noise level: a larger value injects more noise, thereby pushing the initial latent $z_{t}$ closer to the pure-noise prior.
+Specifically, when $\mathrm{DS}=1.0$, the diffusion process starts entirely from noise and discards all information from the inputs, achieving complete object removal but often introducing noticeable global color shifts.
+In contrast, using a slightly lower denoising strength (e.g., $\mathrm{DS}=0.99$) preserves color consistency but risks incomplete object removal or hallucinated unwanted objects.
+Motivated by these observations, we propose Spatially-Varying Denoising Strength (SVDS), which applies $\mathrm{DS}=1.0$ within the masked object region and $\mathrm{DS}=0.99$ in the unmasked background, via re-injecting the background during inference.
+This strategy achieves complete object-effect removal while maintaining background color consistency and preventing edge artifacts during Attention-Guided Fusion (see ablation in the supp.).
 4 Experiments
@@
 The best and second performances are marked in red and orange, respectively.
-Although ObjectClear takes only the object mask as input, it outperforms previous methods, even when those methods are provided with masks that cover both the object and its associated effect regions.
+Although ObjectClear takes only the object mask as input, it outperforms prior methods on most metrics, even when some methods are provided with masks that cover both the object and its associated effect regions.
 Mask Datasets RORD-Val OBER-Test OBER-Wild
 […]
-Object-Effect Mask SDXL-INP [1] 19.39 22.98 0.2432 0.1024 24.07 26.93 0.1296 0.0579 0.8356
-PowerPaint [43] 19.87 21.98 0.2303 0.0776 26.20 27.41 0.1243 0.0409 0.8878
-BrushNet [12] 16.82 18.87 0.3434 0.1692 20.96 23.69 0.2052 0.1180 0.6415
-DesignEdit [10] 20.69 22.49 0.2946 0.1200 26.59 27.63 0.1777 0.0629 0.9150
-CLIPAway [7] 18.87 20.78 0.3328 0.0969 25.38 26.28 0.1039 0.0349 0.8646
-FreeCompose [5] 19.67 21.72 0.3316 0.1110 23.39 25.15 0.1305 0.0629 0.8976
-Attentive Eraser [31] 20.33 21.98 0.2545 0.1015 27.42 29.27 0.1114 0.0249 0.9095
-RORem [15] 21.61 23.11 0.3224 0.0767 27.23 27.95 0.1042 0.0234 0.9157
-Object Mask SDXL-INP [1] 20.23 24.83 0.2042 0.0868 22.42 25.77 0.1428 0.0771 0.8044
-PowerPaint [43] 21.46 24.62 0.1801 0.0648 22.76 24.67 0.1544 0.0729 0.8538
-BrushNet [12] 18.06 23.44 0.2757 0.1821 21.19 24.38 0.1822 0.1123 0.7296
-DesignEdit [7] 22.09 24.26 0.2501 0.1021 24.63 25.48 0.1870 0.0788 0.9002
-CLIPAway [10] 20.58 23.21 0.2770 0.0785 22.32 24.05 0.1357 0.0765 0.8646
-FreeCompose [5] 20.39 22.91 0.3015 0.0897 22.77 24.46 0.1393 0.0690 0.8688
-Attentive Eraser [31] 22.17 24.59 0.1883 0.0643 25.70 27.08 0.1201 0.0437 0.9008
-RORem [15] 22.49 24.10 0.2943 0.0634 24.51 25.28 0.1288 0.0460 0.9025
-ObjectClear (Ours) 26.40 30.75 0.0802 0.0350 32.98 35.14 0.0352 0.0099 0.9172
+Object-Effect Mask SDXL-INP [26] 19.39 22.98 0.2432 0.1024 24.07 26.93 0.1296 0.0579 0.6983
+PowerPaint [48] 19.87 21.98 0.2303 0.0776 26.20 27.41 0.1243 0.0409 0.8044
+BrushNet [12] 16.82 18.87 0.3434 0.1692 20.96 23.69 0.2052 0.1180 0.5358
+DesignEdit [10] 20.69 22.49 0.2946 0.1200 26.59 27.63 0.1777 0.0629 0.8215
+CLIPAway [6] 18.87 20.78 0.3328 0.0969 25.38 26.28 0.1039 0.0349 0.7705
+FreeCompose [4] 19.67 21.72 0.3316 0.1110 23.39 25.15 0.1305 0.0629 0.7555
+Attentive Eraser [34] 20.33 21.98 0.2545 0.1015 27.42 29.27 0.1114 0.0249 0.7940
+RORem [18] 21.61 23.11 0.3224 0.0767 27.23 27.95 0.1042 0.0234 0.8164
+Object Mask SDXL-INP [26] 20.23 24.83 0.2042 0.0868 22.42 25.77 0.1428 0.0771 0.6971
+PowerPaint [48] 21.46 24.62 0.1801 0.0648 22.76 24.67 0.1544 0.0729 0.7699
+BrushNet [12] 18.06 23.44 0.2757 0.1821 21.19 24.38 0.1822 0.1123 0.6341
+DesignEdit [10] 22.09 24.26 0.2501 0.1021 24.63 25.48 0.1870 0.0788 0.8163
+CLIPAway [6] 20.58 23.21 0.2770 0.0785 22.32 24.05 0.1357 0.0765 0.7705
+FreeCompose [4] 20.39 22.91 0.3015 0.0897 22.77 24.46 0.1393 0.0690 0.7451
+Attentive Eraser [34] 22.17 24.59 0.1883 0.0643 25.70 27.08 0.1201 0.0437 0.8080
+RORem [18] 22.49 24.10 0.2943 0.0634 24.51 25.28 0.1288 0.0460 0.8121
+OmniEraser [37] 21.79 22.98 0.2195 0.0542 24.44 24.87 0.1783 0.0142 0.7655
+GeoRemover [47] 24.38 26.47 0.1924 0.0575 24.92 25.33 0.1976 0.0670 0.7938
+OmniPaint [42] 22.75 24.66 0.1178 0.0354 29.06 30.04 0.0521 0.0152 0.8656
+ObjectClear (Ours) 26.24 29.78 0.1157 0.0299 33.04 35.62 0.0342 0.0103 0.8470
+Figure 4: Object Removal on OBER-Test and RORD-Val. Our ObjectClear effectively removes both the masked objects and their associated effects, including shadows and mirror reflections.
 Implementation.
-Our proposed method is built on SDXL-Inpainting [1] and fine-tuned with our OBER dataset with an input resolution of $512\times 512$.
+Our network is built on SDXL-Inpainting [26] and fine-tuned with our OBER dataset with an input resolution of $512\times 512$.
 Training is conducted with a total batch size of $32$ on 8 A100 GPUs for 100k iterations and a learning rate of 1e-5.
-All experiment results are attained with a guidance scale of $7.5$ and $20$ denoising steps.
+All experiment results are attained with a guidance scale of $1.0$ and $20$ denoising steps.
+Figure 5: Object Removal on OBER-Wild.
+Prior methods may not completely remove object effects and can also introduce residual artifacts or unwanted content.
+In contrast, ObjectClear achieves precise object–effect removal while preserving high background fidelity.
 Evaluation Data.
-We evaluate the object removal performance of our method, ObjectClear, against existing approaches on three test datasets: (1) RORD-Val: RORD [29] is a widely used object removal dataset.
-Each sample contains an image pair (images with and without the target object) and a coarse mask covering both the target object and its associated effect.
+We evaluate ObjectClear and prior methods on three test sets: (1) RORD-Val.
+RORD [31] is a widely used object removal dataset.
+Each sample contains an image pair (w/ and w/o object) and a coarse mask covering both object and its effect.
 To avoid duplicated scenarios, we randomly select one image per scene, resulting in a RORD-Val with 343 samples.
-Since most of the object removal methods, including ObjectClear, require only the object mask as input, we augment RORD-Val with accurate object masks manually to enable a more comprehensive evaluation.
-(2) OBER-Test: As described in Sec.
-3.1, we split our collected OBER dataset into a training set and a test set.
-The test set contains 160 samples, each comprising image pairs along with precise masks for both the object and its effect.
-(3) OBER-Wild: To further assess the performance of different methods on out-of-distribution scenarios in the wild, we collect 300 high-quality images featuring objects with associated effects (shadows or reflections) from the Internet.
-We annotate the object masks using DINO [41] and SAM [13], and manually annotate the effect masks.
-Note that the OBER-Wild set does not include the removal ground truth.
-Figure 4: Object Removal on OBER-Test and RORD-Val. Our ObjectClear effectively removes both the masked objects and their associated effects.
-Figure 5: Object Removal on OBER-Wild.
-Our ObjectClear not only effectively removes single objects with their shadows and reflections (top five samples), but also accurately removes the target object when multiple mutually occluding objects exist (bottom two samples).
+Since some object removal methods, including OmniEraser [37], OmniPaint [42] and ours, require only the object mask as input, we manually refine the object masks to ensure fair evaluation.
+(2) OBER-Test.
+As described in Sec.
+3.1, we split our OBER dataset into training and test sets.
+The test set contains 163 samples, each providing image pairs and precise masks for both object and effect.
+(3) OBER-Wild.
+To assess generalization in real-world scenarios, we collect 302 in-the-wild images featuring objects with shadows or reflections.
+Object masks are annotated using DINO [44] and SAM [13], while effect masks are manually labeled.
+Ground-truth removal images are not available for this set.
 Evaluation Metrics.
 For RORD-Val and OBER-Test, where ground truth is available, we evaluate performance using fidelity metrics: PSNR and PSNR-BG (computed only on background regions), as well as two perceptual metrics: LPIPS [45] and CLIP [27] (feature distance).
-For OBER-Wild, which lacks ground truth, we employ ReMOVE [3] to measure the consistency between the object removal region and the surrounding background.
-Figure 6: Object-Effect Attention Map.
-ObjectClear achieves a relatively high recall in the object-effect attention map, with recall values increasing as the denoising step progresses.
-The attention map obtained in the final step effectively covers both the target object and its associated effects, including the object’s shadow (red box) and its reflection in the mirror (green box).
+For OBER-Wild, which lacks ground truth, we evaluate results using ReMOVE, a modified version of ReMOVE [2] designed to better assess visual harmony.
+Specifically, ReMOVE measures the consistency between the output’s removal region and the input background, instead of comparing against the output background as in the original ReMOVE.
 4.1 Comparisons with State-of-the-Art Methods
-We compare ObjectClear with open-source state-of-the-art methods in both image inpainting and object removal.
-The inpainting methods include SDXL-INP [1], PowerPaint [43], and BrushNet [12].
-The object removal methods include CLIP-Away [7], DesignEdit [10], RORem [15], FreeCompose [5], and Attentive Eraser [31].
-Following MULAN [32], we use the text prompt “an empty scene” for the text-guided baselines when necessary.
+We compare ObjectClear with state-of-the-art methods, including: 1) image inpainting methods of SDXL-INP [26], PowerPaint [48], and BrushNet [12]; 2) object removal methods of CLIP-Away [6], DesignEdit [10], RORem [18], FreeCompose [4], Attentive Eraser [34], OmniEraser [37], GeoRemover [47], and OmniPaint [42].
 Quantitative Evaluation.
 Unlike our ObjectClear, which explicitly handles both the target object and its associated effects, most baseline methods operate only within the masked regions, overlooking surrounding areas that are visually affected by the object but not explicitly included in the mask.
-To ensure a fair and comprehensive comparison, we evaluate all methods under two mask conditions: (1) the input mask covers only the target object, and (2) the input mask covers both the target object and its visual effects.
-These are referred as Object Mask and Object-Effect Mask settings in Table 1, respectively.
-As shown in Table 1, ObjectClear achieves state-of-the-art performance across all test sets and metrics.
+To ensure a fair and comprehensive comparison, we test all methods under two mask conditions: (1) Object Mask, and (2) Object-Effect Mask, as reported in Table 1.
+ObjectClear achieves state-of-the-art performance on nearly all test sets and metrics.
 Notably, even when using only the object mask, ObjectClear surpasses methods that rely on both object and effect masks.
 In particular, it achieves a significant advantage in the PSNR-BG metric, highlighting its superior ability to preserve background consistency with the input.
+Table 2: Quantitative Results of Ablation Studies.
+Based on our camera-captured data ($C$ Data), object removal performance improves progressively by introducing the Adaptive Target-Aware Attention (ATA), simulated training data (Sim.
+Data), Attention-Guided Fusion (AGF), and Spatially-Varying Denoising Strength (SVDS) strategy.
+[…]
+(a) 0.99 27.29 27.96 0.0910 0.0247
+(b) 0.99 27.56 28.37 0.0845 0.0217
+(c) 0.99 28.04 28.80 0.0805 0.0196
+(d) 0.99 32.77 35.50 0.0348 0.0106
+(e) 1.00 31.49 33.46 0.0375 0.0120
+(f) SVDS 33.04 35.62 0.0342 0.0103
+Figure 6: Object-Effect Map from Adaptive Target-Aware Attention.
+ObjectClear achieves a relatively high recall in the object-effect map, with recall values increasing as the denoising step progresses.
+The attention map obtained in the final step effectively covers both the target object and its associated effects, including the object’s shadow (red box) and its reflection in the mirror (green box).
+Figure 7: Effectiveness of Attention-Guided Fusion.
+The difference maps visualize the pixel-wise differences between input and output.
+RORem [18] fails to remove the reflection and alters the background, yielding low differences in the reflection area and high values in the background.
+In contrast, both ObjectDrop [38] and our ObjectClear successfully remove the reflection, with ObjectClear better preserving the background as indicated by lower differences.
 Qualitative Evaluation.
@@
 5.
-Generation-based inpainting approaches, such as SDXL-INP [24], BrushNet [12], and PowerPaint [43], often generate new objects within the masked regions while fail to remove the effects of the removed objects.
-In contrast, previous object removal methods, such as Attentive Eraser [31], DesignEdit [10], and RORem [15] demonstrate strong performance in removing the target object itself, but stil fail to eliminate the associated effects.
-Thanks to our carefully curated training data and the proposed object-effect attention mechanism, ObjectClear effectively removes not only the target object but also its associated shadows and reflections.
-Moreover, as shown in the last two rows of Fig.
-5, ObjectClear effectively handles complex cases involving multiple mutually occluding objects.
-Table 2: Quantitative Results of Ablation Studies.
-Based on our camera-captured data (CC Data), our object removal performance improves progressively by adding the object-effect mask loss ($\mathcal{L}_{mask}$) and simulated training data (Sim.
-Data).
-Further improvement is achieved through the proposed Attention-Guided Fusion (AG Fusion).
-[…]
-(a) 27.29 27.96 0.0910 0.0247
-(b) 27.56 28.37 0.0845 0.0217
-(c) 28.00 28.81 0.0818 0.0190
-(d) 32.98 35.14 0.0352 0.0099
+Generation-based inpainting approaches, such as SDXL-INP [26], and PowerPaint [48], often generate new objects within the masked regions while fail to remove the effects of the removed objects.
+In contrast, previous object removal methods, such as Attentive Eraser [34], DesignEdit [10], and RORem [18] demonstrate strong performance in removing the target object itself, but still fail to eliminate the associated effects.
+OmniEraser [37] demonstrates the ability to remove both objects and their effects, though it may still leave residual objects or effects, occasionally erase non-target regions, or introduce background changes.
+Similarly, OmniPaint [42] maintains good background fidelity but may occasionally fail to fully remove the target object and its associated effects.
+Benefiting from our training data and network designs, ObjectClear excels at accurately removing both the target object and its effects.
+Moreover, it robustly handles complex scenarios involving multiple occluding objects (see the last row of Fig.
+5).
 4.2 Ablation Studies
-The Effectiveness of the Object-Effect Mask Loss $\mathcal{L}_{mask}$.
-The object-effect mask loss is designed to explicitly guide ObjectClear to attend to both the target object and its associated effect regions, leading to more thorough and natural removal results.
-As shown in the chart on the left of Fig.
-6, the predicted object-effect attention map achieves a high recall of 0.97.
-Besides, the attention map in the last column of Fig.
-6 on the right, which is from the final denoising step (Step 20), effectively reveals not only the object’s shadow (marked with a red box), but also the object’s reflection in the mirror (marked with a green box), altogether leading to complete and visually coherent object removal results.
-From Table 2 (a) to (b), improvements can be seen across all metrics with $\mathcal{L}_{mask}$ added, indicating its effectiveness in improving object removal accuracy.
+The Effectiveness of Adaptive Target-Aware Attention (ATA).
+The ATA is designed to explicitly guide ObjectClear to attend to both the target object and its effects, leading to more complete and precise removal.
+From Table 2 (a) to (b), improvements can be seen across all metrics with ATA introduced, indicating its effectiveness in improving object removal accuracy (see visual comparison in supp).
+Besides, as shown in the plot of Fig.
+6 (left), the object-effect map from ATA module achieves a high recall of 0.97.
+In Fig.
+6 (right), the attention map from the final denoising step (Step 20) effectively reveals object’s shadow (red box) and mirror reflection (green box), altogether leading to complete and visually coherent object removal results.
 The Effectiveness of Simulated Data.
-The simulated data effectively enhances the background diversity and scenario complexity of the training set, leading to notable improvements in object removal performance and better background preserving, as shown in Table 2(c).
-Our carefully designed simulation pipeline, especially the multi-object data synthesis, enables ObjectClear to handle challenging cases involving multiple mutually occluding objects, as demonstrated in the last two rows of Fig.
-5.
-More visualized results can be found in the appendix.
-Effectiveness of the Attention-Guided Fusion.
-The attention-guided fusion operation is designed to preserve background consistency during object removal, a common challenge in generation-based methods such as ObjectDrop [34] and RORem [15], as illustrated in Fig.
+The simulated data effectively enhances the scale and diversity of the training data, leading to notable improvements in object removal performance and better background preserving, as shown in Table 2(c).
+Our carefully designed simulation pipeline, especially the multi-object data synthesis, further enables ObjectClear to tackle challenging scenarios involving multiple objects.
+We provide more visual results in supplementary.
+Effectiveness of the Attention-Guided Fusion (AGF).
+The AGF is designed to preserve background consistency after object removal, a common challenge in generation-based methods like ObjectDrop [38] and RORem [18], as illustrated in Fig.
 7.
-As the denoising process progresses, the proposed object-effect attention yields increasingly accurate attention maps (Fig.6), which we leverage to guide the fusion of background images.
-This strategy significantly improves background preservation, as evidenced by notable gains in PSNR-BG.
-Consequently, we observe a substantial boost in overall performance, as reflected in the quantitative results from (c) to (d) in Table 2, where PSNR improves markedly from $28.00$ to $32.98$.
-Figure 7: Effectiveness of Attention-Guided Fusion.
-The difference maps visualize the pixel-wise differences between the input and output images.
-RORem [15] fails to remove the reflection of the balloon and introduces changes to the background, resulting in relatively high difference values in the background area and low values in the reflection region.
-In contrast, both ObjectDrop [34] and our ObjectClear successfully remove the reflection.
-Notably, ObjectClear better preserves the background, thanks to our predicted object-effect map and proposed attention-guided fusion strategy.
-4.3 More Applications
-ObjectClear can be flexibly extended to various applications, such as object insertion and movement.
-We also utilize our OBER dataset to train an insertion network that is able to generate natural object effects.
-Please refer to the appendix for more details.
-As shown in Fig.
-8, even when only the target objects are specified for insertion and movement, ObjectClear is capable of generating plausible and natural shadows and reflections accordingly.
-Figure 8: Object Insertion and Movement. In addition to accurately inserting or repositioning objects, our ObjectClear also generates plausible and natural shadows and reflections accordingly.
+As the denoising process, our Adaptive Target-Aware Attention yields increasingly accurate attention maps (Fig.6), which we leverage to guide the fusion of input background.
+This strategy significantly improves background preservation, as evidenced by large gains in PSNR-BG.
+Consequently, we observe a marked boost in overall performance, with PSNR increasing from $28.04$ to $32.77$, as shown in Table 2(c–d).
+Effectiveness of the Spatial-Varying Denoising Strength.
+The proposed Spatially-Varying Denoising Strength (SVDS) addresses two major issues: incomplete object removal or hallucinated objects when using $\mathrm{DS}=0.99$, and global color shifts when using $\mathrm{DS}=1.0$ (see visual ablation in supp.).
+As shown in Table 2(d–f), SVDS (f) outperforms both $\mathrm{DS}=0.99$ (d) and $\mathrm{DS}=1.0$ (e).
+It effectively facilitates complete object removal while maintaining consistent background colors and details.
 5 Conclusion
-We introduce ObjectClear, a practical framework for object removal that ensures high-quality object-effect removal and background-preserving results across diverse real-world scenarios.
-Our method employs an Object-Effect Attention (OEA) mechanism to adaptively focus on removal regions, including both the object and its effects, coupled with Attention-Guided Fusion (AGF) to preserve background details accurately.
-We also propose OBER, a large-scale and diverse hybrid dataset combining camera-captured and simulated data.
-With the carefully designed mechanism and curated dataset, ObjectClear achieves robust and superior performance, effectively addressing existing challenges in object removal tasks.
-In addition, ObjectClear can be easily adapted to various applications, such as object insertion and object movement, and consistently achieves superior performance.
-These advancements position ObjectClear as a reliable solution for practical image editing, laying a solid foundation for future research in controllable image processing through local editable regions.
+We introduce ObjectClear, a practical framework for object removal that achieves high-quality object-effect removal while maintaining background consistency across diverse real-world scenarios.
+The framework employs an Adaptive Target-Aware Attention to adaptively focus on removal regions, along with Attention-Guided Fusion and Spatial-Varying Denoising Strength strategy to preserve background details.
+In addition, we present OBER, a large-scale and diverse hybrid dataset that integrates camera-captured and simulated data.
+Thanks to our dataset and network designs, ObjectClear shows robust and superior performance, effectively overcoming key challenges in object removal and laying a solid foundation for future research in this field.
+Acknowledgement. This research is supported by cash and in-kind funding from NTU S-Lab and industry partner(s).
 Appendix
-In this supplementary material, we provide additional discussions and results to complete the main paper.
+In this appendix, we provide additional discussions and results to complete the main paper.
 In Sec.
-A, we present further details on data augmentation during training and other extended applications.
+6, we present further implementation details on data augmentation during training and other extended applications.
 In Sec.
-B, we provide more details of the proposed OBER dataset, including statistics and some examples for demonstration.
+7, we provide a more detailed introduction to the proposed OBER dataset, including statistics and some examples for demonstration.
 In Sec.
-C, we report additional results, such as more results on ablation studies, a user study, experiments with user strokes, more results of object insertion and movement, and additional comparisons on in-the-wild data.
-A More Details of ObjectClear
-A.1 Training Augmentations
+8, we report additional results, such as further ablations, a user study, experiments with user strokes, results for object insertion and movement, and additional comparisons on in-the-wild images.
+In addition, we provide an online [interactive demo] on Hugging Face, enabling users to remove arbitrary objects through simple clicks.
+6 Implementation Details
+6.1 Training Augmentations
 During training, we apply on-the-fly random cropping to enable the network to learn object removal across varying object sizes.
@@
 As demonstrated in the qualitative results in Sec.
-C.3, our method effectively handles coarse user-drawn masks by implicitly completing and refining them, showcasing strong robustness to imprecise inputs.
-A.2 More Details for Other Applications
-In this section, we further discuss implementation details of other applications, including object insertion and object movement.
+8.6, our method effectively handles coarse user-drawn masks by implicitly completing and refining them, showcasing strong robustness to imprecise inputs.
+6.2 More Details for Extended Applications
+ObjectClear can be flexibly extended to various applications.
+Visual results of object insertion and movement are shown in Fig.
+16, where our models generate realistic visual effects.
+In this section, we provide the implementation details of object insertion and object movement.
 Object Insertion.
-The insertion network also leverages the OBER dataset for training and adopts the same architecture as ObjectClear (Fig.
-3 of the main paper), thus receiving the input tuple of $<z_{t},I_{in},M_{o},c>$ and supervised by $I_{GT}$ (for output image) and $M_{fg}$ (for object-effect attention).
-However, $I_{in}$, $I_{GT}$, and $c$ are constructed in an “reverse” way compared with the removal network.
-While the ground truth $I_{GT}$ is the original image with the object accompanied by natural effects, $I_{in}$ is the one with the object simply copied and pasted on the background.
-Specifically, to generate $I_{in}$, we first extract the object from $I_{GT}$ with the object mask $M_{o}$ only, and then we paste it onto the corresponding background image.
+The insertion network also leverages the OBER dataset for training and adopts the same architecture as ObjectClear, as illustrated in Fig.3 of the main paper, which receives the input tuple of $z_{t},I_{in},M_{o},c>$ and supervised by $I_{GT}$ (for output image) and $M_{fg}$ (for Adaptive Target-Aware Attention).
+However, $I_{in}$, $I_{GT}$, and $c$ are constructed in a reverse manner compared with the removal network.
+Specifically, the ground-truth image $I_{GT}$ corresponds to the original image containing the object along with its associated effects, while $I_{in}$ is obtained by simply copying and pasting the object onto the background without any effects.
+To generate $I_{in}$, we first extract the object from $I_{GT}$ using the object mask $M_{o}$ only, and then paste it into the corresponding background image.
 In addition to the image pair, the input text $c$ is also reversed to “insert the instance of”.
-Notably, the object-effect attention map is still supervised by $M_{fg}$.
-In this context, $M_{fg}$ refers to the mask that covers both the object and its generated effects (e.g., shadows or reflections).
-During inference, we directly paste an object (without effects) onto the background scene, along with its object mask $M_{o}$, which forms the input image $I_{in}$.
+Notably, the Adaptive Target-Aware Attention map is still supervised by $M_{fg}$.
+Here, $M_{fg}$ refers to the mask that covers both the object and its generated effects (e.g., shadows or reflections).
+During inference, we obtain the input image $I_{in}$ by pasting an object foreground (w/o effects) onto the background image, and we feed it together with its corresponding object mask $M_{o}$ as the input pair.
 The network then generates the output image where both the object and its generated effects are harmoniously inserted.
-Since the insertion network also integrates the object-effect attention, we can also apply Attention-Guided Fusion to preserve the background fidelity while generating realistic object effects.
+Since the insertion network also integrates the Adaptive Target-Aware Attention, we further apply the Attention-Guided Fusion to preserve background fidelity while synthesizing realistic object effects.
 Object Movement.
-To enable object movement, we combine the ObjectClear and insertion network introduced above.
-Specifically, we first apply ObjectClear to remove the target object along with its associated effects (e.g., shadows or reflections), resulting in a clean object-free background.
-The object is then extracted using its given object mask.
-Users are allowed to specify a new location and optionally adjust the object scale before the object is re-inserted into the clean background.
+To enable object movement, we combine our object removal (ObjectClear) and insertion models into a two-stage framework.
+Specifically, we first apply ObjectClear to remove the target object and its associated effects, producing a clean, object-free background.
+The object is then extracted using its provided object mask, and users can specify a new location and adjust its scale before reinsertion.
 With our insertion network, the object is harmonized with the new context by generating realistic effects.
-This two-stage approach supports controllable object movement while ensuring visual realism and consistency.
-B More Details of OBER Dataset
+Without retraining any additional networks, this two-stage pipeline supports controllable object movement while maintaining visual realism and consistency.
+7 More Details of OBER Dataset
+Figure 8: Samples from our OBER Dataset.
+We provide some samples from our OBER dataset including both camera-captured data and simulated data.
+While all data from both categories are annotated with fine-grained object masks and object-effect masks, we also extract RGBA foreground objects with associated effects from camera-captured data (as discussed in Sec.
+3.1 in the main paper), which could be used to construct realistic simulated samples.
+In the simulated samples, we not only include the shadow effects but also the reflection effects, which are rarely included in other datasets, thanks to our reflection pair simulation strategy detailed in Sec.
+7.
 Table 3: Dataset Overview.
@@
 Camera-Captured Simulated OBER-Test OBER-Wild
-#Image Pairs 2,715 10,000 160 300
+#Image Pairs 2,715 10,000 163 302
 w/ Ground Truth ×
@@
 w/ Object-Effect Mask
-Table 4: Comparison of OBER Dataset with Existing Datasets. * indicates that the Counterfactual dataset is not publicly available.
-Description RORD [29] MULAN [32] DESOBAv2 [18] Counterfactual [34] OBER (ours)
+Overview.
+Table 3 provides an overview of our OBER dataset [link].
+The training set consists of 2,715 camera-captured image pairs and 10,000 simulated pairs, all annotated with object masks and object-effect masks.
+For evaluation, we provide two test subsets: OBER-Test (includes 163 pairs with ground truth) and OBER-Wild (includes 302 pairs without ground truth), where object masks and object-effect masks are also available.
+Figure 8 presents samples from our OBER dataset, including camera-captured and simulated data with annotations such as object masks, object-effect masks, and RGBA foregrounds.
+Reflection Pair Simulation.
+Thanks to the strong priors of generative models, we observed that a model trained with only limited indoor mirror-reflection data can still generalize to removing simple outdoor reflection effects (e.g., water reflections), which is consistent with the findings in ObjectDrop [38].
+However, the model often fails in more challenging cases, particularly when the object and its reflection are spatially separated or when the reflection is heavily distorted by water ripples.
+To address this limitation, we adopt a human-in-the-loop strategy to collect paired reflection data.
+Specifically, we first use our trained model to perform inference on 200 real-world reflection images, and then manually select 50 high-quality results (see examples in Fig.
+8(b)), which are then added as an important supplement to the training data.
+We found that even a small amount of high-quality reflection pairs can significantly improve the model’s generalization ability across diverse reflection scenarios.
+Comparison with Existing Datasets.
+We compare our OBER dataset with existing datasets, including those focused on shadow removal (DESOBA-v2 [20]) and object removal (RORD [31], MULAN [35], Counterfactual Dataset [38], Video4Removal [37]).
+As summarized in Table 4, unlike prior datasets, OBER provides all three types of mask annotations as well as RGBA object foregrounds, enabling a wide range of tasks, including object removal, effect removal, and joint object–effect removal.
+DESOBA-v2 [20] targets only effect removal and thus does not handle the objects.
+In contrast, MULAN [35] focuses solely on object removal without associated effects such as shadows or reflections.
+Moreover, the ground-truth images in both DESOBA-v2 and MULAN are produced by existing inpainting models, which may limit their visual realism.
+RORD [31] is among the first to provide object–effect masks, however, these masks are very coarse and can not separately annotate objects and their effects.
+Consequently, it cannot support independent removal of either the object or the effect.
+The Counterfactual Dataset [38] and Video4Removal Dataset [37] are designed for object–effect removal tasks, the same with ours, but they are not publicly accessible.
+Furthermore, these datasets provide only object masks, lacking effect masks and RGBA foregrounds, which limits their scalability and diverse usage.
+In contrast, our OBER dataset provides richer annotations, including precise and separate masks for objects and their effects, together with RGBA foregrounds, enabling more flexible and fine-grained removal tasks.
+Notably, the effect masks serve as crucial supervision for learning accurate object–effect removal while preserving background fidelity.
+OBER is a hybrid dataset comprising high-quality captured and realistic synthetic data, covering diverse and complex scenarios such as multi-object occlusions and indoor/outdoor reflections.
+We will release our OBER dataset publicly, which we believe will significantly benefit future research in this field.
+Table 4: Comparison of OBER Dataset with Existing Datasets. * indicates that the dataset is not publicly available.
+Description RORD [31] MULAN [35] DESOBAv2 [20] Counterfactual [38] Video4Removal [37] OBER (ours)
 Tasks Object Removal × ×
@@
 Annotations Object Mask ×
-Effect Mask × × ×
-Object-Effect Mask ✓(Coarse) × ×
-RGBA Objects × × × ×
+Effect Mask × × × ×
+Object-Effect Mask ✓(Coarse) × × ×
+RGBA Objects × × × × ×
 Multi Objects ×
-Effect Types Shadow × Shadow Shadow Shadow & Reflection
 Camera-Captured GT × ×
-Overview.
-Table 3 provides an overview of our OBER dataset.
-The training set consists of 2,715 camera-captured image pairs and 10,000 simulated pairs, all annotated with object masks and object-effect masks.
-For evaluation, we provide two test subsets: OBER-Test (includes 160 pairs with ground truth) and OBER-Wild (includes 300 pairs without ground truth), where object masks and object-effect masks are also available.
-In addition, we showcase some samples from our OBER dataset in Fig.
-9, including camera-captured data and simulated data, alongside with their annotations such as object masks, object-effect masks, and RGBA object foregrounds.
-Reflection Pair Simulation.
-Thanks to the strong priors of generative models, we observed that the model trained without reflection data was still able to remove some simple reflection effects, a finding consistent with the conclusions drawn in ObjectDrop [34].
-However, the model often struggled with more challenging reflection removal cases, such as when the object and its reflection were spatially separated or when the reflection was heavily distorted by surface ripples.
-To address this limitation, we adopted a human-in-the-loop strategy to collect paired reflection data.
-Specifically, we used our trained model to perform inference on 200 real-world reflection images and manually selected 50 high-quality results (as shown in Fig.
-9(b)), which were then added as an important supplement to the training data.
-We found that even a small amount of high-quality reflection data pairs could considerably improve the model’s ability to generalize across diverse reflection scenarios.
-Comparison with Existing Datasets.
-We compare our OBER dataset with existing datasets, including those focused on object removal (RORD [29], MULAN [32], Counterfactual Dataset [34]) and one on shadow removal (DESOBA-v2 [18]).
-As summarized in Table 4, our OBER is the only dataset that includes all three types of mask annotations and RGBA foreground objects, enabling a wide range of tasks such as object removal, effect removal, and joint object-effect removal.
-While RORD [29] is among the first to provide object-effect masks, those masks are coarse without separate annotations for objects and their effects.
-As a result, it cannot support independent removal of either the object or the effect, but only their joint removal.
-In contrast, our OBER dataset provides precise and separate masks for objects and their effects, enabling more flexible and fine-grained removal tasks.
-In addition, MULAN [32] focuses solely on object removal without addressing associated effects such as shadows or reflections.
-DESOBA-v2 [18] targets only effect removal and thus does not handle the objects.
-Moreover, the ground truth images in both datasets are not captured by cameras but are instead synthesized using generative models, which may limit their realism.
-The Counterfactual Dataset proposed in ObjectDrop [34] is not publicly available, and it only includes object masks without object-effect masks.
-It also lacks foreground RGBA data for objects and associated effects, limiting its extensibility and applicability to more advanced object editing tasks.
-Figure 9: Samples from our OBER Dataset.
-We provide some samples from our OBER dataset including both camera-captured data and simulated data.
-While all data from both categories are annotated with fine-grained object masks and object-effect masks, we also extract RGBA foreground objects with associated effects from camera-captured data (as discussed in Sec.
-3.1 in the main paper), which could be used to construct realistic simulated samples.
-In the simulated samples, we not only include the shadow effects but also the reflection effects, which are rarely included in other datasets (Table 4), thanks to our reflection pair simulation strategy detailed in Sec.
-B.
-C More Results
-C.1 Results for Ablation Study
-Effectiveness of $\mathcal{L}_{mask}$.
-Thanks to the rich annotations in our OBER dataset, the object-effect mask enables a supervision loss, denoted as $\mathcal{L}_{mask}$, which guides the cross-attention layers to focus on both the object and its associated effects, while preserving background textures.
-This supervision facilitates a decoupled optimization of object removal and background reconstruction.
+8 More Results
+8.1 Results for Ablation Study
+Effectiveness of Adaptive Target-Aware Attention.
+Thanks to the OBER dataset’s rich annotations, Adaptive Target-Aware Attention (ATA) leverages the object-effect mask and its supervision loss $\mathcal{L}_{mask}$.
+It guides cross-attention layers to focus on the object and its associated effects while preserving background textures, which enables decoupled optimization of object removal and background reconstruction.
 As shown in Fig.
-10, with the mask loss $\mathcal{L}_{mask}$, the network can adaptively identify the object-effect regions to be removed, as reflected in the attention mask (shown in yellow boxes).
+9, ATA adaptively identifies object-effect regions to be removed, as reflected in the attention mask (shown in yellow boxes).
 This leads to more accurate and complete removal of objects and their shadow effects, without mistakenly erasing unrelated background content.
-Figure 10: Effectiveness of $\mathcal{L}_{mask}$.
-It could be observed that when training without the $\mathcal{L}_{mask}$ supervision, the model struggles to accurately remove the target object and its effects, leading to mistakenly erasing unrelated background object (right) or effects (left).
-In contrast, when supervised with $\mathcal{L}_{mask}$, the attention mask (shown in yellow boxes) could identify the removal regions well, thus leading to more accurate and complete removal results.
-Effectiveness of Simulated Data.
-To balance data realism and scalability, in addition to the camera-captured data, we scale up our OBER dataset with a carefully designed simulation pipeline.
-Our simulation data is generated by compositing the foreground RGBA object (extracted from camera-captured data) onto diverse backgrounds.
-In particular, the simulation of multi-object compositions leads to notable improvements in object removal robustness when mutually occluding objects exist, as shown in Fig.
-11 (left).
-Furthermore, our simulated reflection image pairs (discussed in Sec B) greatly improve the model capability of removing reflections, including challenging cases with significant distortion due to water surface ripples, as shown in Fig.
-11 (right).
+Figure 9: Effectiveness of Adaptive Target-Aware Attention (ATA).
+It could be observed that when training without ATA (w/o $\mathcal{L}_{mask}$), the model struggles to accurately remove the target object and its effects, leading to mistakenly erasing unrelated background object (right) or effects (left).
+In contrast, when introducing TAT trained with the mask supervision $\mathcal{L}_{mask}$, the attention maps (shown in yellow boxes) can accurately localize the removal regions, leading to more precise and complete removal results.
 Figure 10: Effectiveness of Simulated Data.
 Since our simulation data includes multi-object compositions, training with such data enables the model to accurately remove the target object and its associated effects while preserving unrelated object effects (left).
-In addition, adding the reflection data pairs during training greatly enhances the model capability of removing reflections, including challenging cases with significant distortion due to water surface ripples (right).
+In addition, adding the reflection data pairs during training greatly enhances the model capability of removing reflections, even in challenging cases (right).
+Figure 11: Effectiveness of Spatially-Varying Denoising Strength (SVDS).
+$\mathrm{DS}=0.99$ often leads to incomplete removal or hallucinated objects, while $\mathrm{DS}=1.0$ causes noticeable color inconsistency (shown after AGF, where the background is from the input image and the object/affected areas are from the removal results).
+In contrast, SVDS achieves complete object removal with consistent background colors.
+Figure 12: Effectiveness of Attention-Guided Fusion (AG Fusion).
+We visualize background detail changes by computing difference maps between the generated image and input.
+Ideally, these maps should exhibit large values only within the object removal regions, covering the object and its effects, while remaining minimal in the unrelated background areas.
+As observed, without AG Fusion, noticeable differences appear in the background.
+In contrast, with the AG Fusion strategy, these undesired background discrepancies are largely eliminated.
+Effectiveness of Simulated Data.
+To balance realism and scalability, in addition to camera-captured data, we augment OBER with a simulation pipeline.
+Our simulation data is generated by compositing the RGBA foreground objects (extracted from camera-captured data) onto diverse backgrounds.
+In particular, the simulation of multi-object compositions leads to notable improvements in object removal robustness under mutual occlusions, as shown in Fig.
+10 (left).
+Furthermore, our simulated reflection pairs greatly improve the model capability of removing reflections, even in challenging cases (Fig.
+10, right).
 Effectiveness of Attention-Guided Fusion.
@@
 12, we visualize the background detail changes by showing the difference maps between the generated image and the corresponding input, where a clear improvement on background preservation could be observed.
-Figure 12: Effectiveness of Attention-Guided Fusion (AG Fusion).
-We visualize the background detail changes by showing the difference maps between the generated image and the corresponding input.
-In ideal cases, the difference maps should show large values only in the object removal regions, including the object and its effects, while showing little difference in the unrelated background.
-It could be observed that when AG Fusion is not employed, although the difference in object removal regions is significant, the difference in the unrelated background is also noticeable.
-However, with the AG Fusion strategy, we almost eliminate the undesirable background difference thanks to the accurate attention map predicted.
-C.2 User Study
+Effectiveness of Spatially-Varying Denoising Strength.
+In diffusion-based image editing, the initial latent for denoising is typically obtained by adding noise to the input image latent.
+The denoising strength (DS), $\mathrm{DS}\in[0,1]$, controls the noise level: a larger value injects more noise, thereby pushing the initial noisy latent closer to the pure-noise prior.
+When $\mathrm{DS}=1.0$, the diffusion process starts entirely from noise, discarding information from the input.
+In this paper, we propose Spatial-Varying Denoising Strength (SVDS), which applies $\mathrm{DS}=1.0$ within the masked object region and $\mathrm{DS}=0.99$ (an empirical setting commonly adopted by previous methods [26]) outside in the unmasked background, ensuring complete object removal while maintaining color consistency.
+As shown in Fig.
+11, setting $\mathrm{DS}=0.99$ often leads to incomplete removal or hallucinated objects, whereas $\mathrm{DS}=1.0$ results in noticeable color inconsistency.
+To highlight this inconsistency, the results of $\mathrm{DS}=1.0$ in Fig.
+11 are obtained after the Attention-Guided Fusion (AGF) operation, where the background is taken from the original input image, while the object and affected areas are taken from the removal results.
+In contrast, our method with SVDS achieves superior performance in both object removal and preservation of background color consistency.
+Table 5: Quantitative comparison on OmniPaint-Bench and OmniEraser-Bench. The best and second performances are marked in red and orange, respectively.
+Methods OmniPaint-Bench OmniEraser-Bench
+[…]
+SDXL-INP [26] 20.51 0.1924 0.0929 0.1100 21.26 0.1968 0.0938 0.0926
+PowerPaint [48] 21.55 0.1882 0.0874 0.0990 22.21 0.2001 0.0911 0.0907
+BrushNet [12] 19.00 0.2733 0.1909 0.1446 20.58 0.2098 0.1182 0.1098
+DesignEdit [10] 24.32 0.1763 0.0769 0.0717 23.87 0.2168 0.1020 0.0755
+CLIPAway [6] 20.07 0.1985 0.0929 0.1129 20.78 0.2035 0.0934 0.0956
+FreeCompose [4] 22.25 0.1631 0.0723 0.0770 22.60 0.1782 0.0733 0.0688
+Attentive Eraser [34] 24.17 0.1523 0.0501 0.0590 24.77 0.1538 0.0463 0.0388
+RORem [18] 23.92 0.1462 0.0478 0.0576 23.70 0.1746 0.0532 0.0416
+OmniEraser [37] 23.02 0.2034 0.0564 0.0640 23.83 0.1766 0.0460 0.0481
+OmniPaint [42] 25.56 0.1001 0.0249 0.0367 25.67 0.1069 0.0213 0.0182
+ObjectClear (Ours) 26.35 0.0950 0.0244 0.0334 27.90 0.0942 0.0230 0.0142
+Table 6: Quantitative comparison on MULAN Dataset. The best and second performances are marked in red and orange, respectively.
+[…]
+SDXL-INP [26] 19.91 0.2494 0.1324 0.1312
+PowerPaint [48] 21.18 0.2449 0.1087 0.0962
+BrushNet [12] 18.22 0.3181 0.2062 0.1893
+DesignEdit [10] 23.26 0.2375 0.1114 0.0725
+CLIPAway [6] 20.08 0.2666 0.1180 0.1152
+FreeCompose [4] 21.30 0.2337 0.0828 0.0703
+Attentive Eraser [34] 23.96 0.1960 0.0551 0.0397
+RORem [18] 23.53 0.2369 0.0571 0.0438
+OmniEraser [37] 21.56 0.2642 0.0728 0.0682
+OmniPaint [42] 22.29 0.1915 0.0650 0.0550
+ObjectClear (Ours) 24.89 0.1586 0.0468 0.0373
+8.2 Comparisons on Additional Benchmarks
+To further evaluate the robustness and generalization ability of our method, we conduct additional evaluations on three benchmarks, i.e., OmniEraser-Bench [37], OmniPaint-Bench [42], MULAN [35].
+OmniEraser-Bench and OmniPaint-Bench Datasets.
+OmniEraser-Bench [37] and OmniPaint-Bench [42] are two recent benchmarks for object–effect removal, which align with our task setting.
+For a fair comparison, all methods use their default input sizes (OmniEraser [37] at 1024, all others at 512), then resize the outputs to the same size (short side 512) for evaluation.
+Table 5 shows our approach outperforms all baselines across metrics on both benchmarks.
+MULAN Dataset.
+Since the ground truth of some samples in MULAN [35] retains shadows or reflections, it is not suitable for evaluating object–effect removal.
+Therefore, we randomly sample 500 “effect-free” image pairs from MULAN for our evaluation.
+As shown in Table 6, our method achieves the best performance across all metrics, even surpassing RORem [18], which is trained on MULAN, demonstrating the strong object removal capability of our approach.
+This also confirm that our model performs robustly on “effect-free” object removal datasets, without introducing negative effects when the objects do not show any shadows or reflections.
+8.3 Fair Comparison with Object-Effect Mask
+Our Attention-Guided Fusion (AGF) module leverages object-effect masks predicted by the proposed Adaptive Target-Aware Attention to blend the original input background back into the generated result.
+Importantly, these masks are predicted by our model rather than taken from annotations, ensuring that we do not use any privileged information unavailable to other approaches.
+Furthermore, existing baseline methods do not have the capability to predict object–effect masks and therefore cannot perform background blending in the same way.
+This makes AGF an integral part of our model design rather than an external post-processing step, and the comparisons in the main paper are therefore fair.
+To further demonstrate that the performance gain is not solely due to the background blending, we perform an additional experiment where all baseline methods are given the ground-truth object-effect mask for blending.
+As shown in Table 7, our method continues to outperform all baselines under this setting, indicating that our superior results primarily come from more effective object–effect generation rather than the availability of blending masks.
+Table 7: Quantitative results on RORD-Val with object-effect masks for blending.
+All baseline methods are equipped with background blending using ground-truth object-effect masks.
+The best and second performances are marked in red and orange, respectively.
+Our method using our predicted object-effect masks achieves the best performance across all metrics.
+[…]
+SDXL-INP [26] (w/b) 21.67 0.1592 0.0688 0.0932
+PowerPaint [48] (w/b) 22.51 0.1472 0.0560 0.0606
+BrushNet [12] (w/b) 18.48 0.2421 0.1572 0.1465
+DesignEdit [10] (w/b) 23.73 0.1580 0.0721 0.0755
+CLIPAway [6] (w/b) 21.96 0.1735 0.0666 0.0734
+FreeCompose [4] (w/b) 23.08 0.1603 0.0829 0.0834
+Attentive Eraser [34] (w/b) 22.90 0.1700 0.0880 0.0854
+RORem [18] (w/b) 25.24 0.1398 0.0387 0.0498
+ObjectClear (Ours) 26.24 0.1157 0.0191 0.0299
+8.4 Generalization to Multi-Object Removal
+The OBER dataset also covers multi-object removal cases, and our model generalizes well to such scenarios (Fig.
+13).
+This reflects the strong generalization capability of a network trained on our dataset.
+Figure 13: Multi-object removal. Our method removes one or multiple objects simultaneously using masks to indicate the targets.
+Figure 14: User Study. Our ObjectClear is preferred by human voters over four representative state-of-the-art methods (PowerPaint [48], RORem [18], OmniEraser [37], and OmniPaint [42].
+8.5 User Study
 To enable a more comprehensive evaluation, we conducted a user study on object removal results for in-the-wild images.
-We compared ObjectClear with two representative state-of-the-art methods: PowerPaint [43] and RORem [15].
-For a fair comparison, we evaluated each baseline method under two settings: (1) conditioned on the object mask and (2) conditioned on the object-effect mask.
-In contrast, all ObjectClear results were generated using only the object mask, and we compared our outputs against both settings of the baselines.
+We compared ObjectClear with four representative state-of-the-art methods: PowerPaint [48], RORem [18], OmniEraser [37], and OmniPaint [42].
+For a fair comparison, we evaluated PowerPaint and RORem under two settings: (1) conditioned on the object mask and (2) conditioned on the object-effect mask, OmniEraser and OmniPaint are conditioned only on the object mask.
+All ObjectClear results were generated using only the object mask, and we compared our outputs against both settings of PowerPaint and RORem, as well as the object-mask setting of OmniEraser and OmniPaint.
 We invite a total of 30 participants for this user study.
-Each volunteer was presented 40 randomly selected image quadruples, consisting of: an input image, two results from a baseline method under different mask conditions, and our result.
+Each volunteer was presented 80 randomly selected image quadruples, consisting of: an input image, two results from a baseline method under different mask conditions, and our result (for OmniEraser and OmniPaint, only the object-mask result was provided, consistent with our setting, thus forming a triple set).
 Participants were asked to select the best removal result based on two criteria: the realism of the object region and the preservation of background details.
-As summarized in Table 5, ObjectClear outperforms the baselines under both mask settings.
-Notably, although baseline methods benefited from access to object-effect masks, our ObjectClear still won more user preference with the object mask only.
-Table 5: User Study.
-Our ObjectClear is preferred by human voters over other two representative state-of-the-art methods (PowerPaint [43] and RORem [15]) under both settings.
-Although the compared methods won more user preference given the object-effect mask instead of the object mask, our ObjectClear still received far more user preference with the object mask only.
-Methods PowerPaint (Object Mask) [43] PowerPaint (Object+Effect Mask) [43] ObjectClear (Ours)
-User Preference 2.68% 7.03% 90.29%
-Methods RORem (Object Mask) [15] RORem (Object+Effect Mask) [15] ObjectClear (Ours)
-User Preference 3.51% 11.80% 84.69%
-C.3 Results with User Strokes
+As summarized in Fig 14, ObjectClear outperforms the baselines under both mask settings.
+Notably, although some baseline methods benefited from access to object-effect masks, our ObjectClear won more user preference with the object mask only.
+8.6 Results with User Strokes
 In practical applications, users often interact with visual systems through imprecise or casually drawn inputs, such as rough scribbles or incomplete masks.
@@
 Therefore, it is essential for a robust object removal network to effectively process such arbitrary mask inputs without relying on carefully crafted annotations.
-Benefiting from our mask augmentation strategy and object-effect attention mechanism, our network demonstrates strong robustness to diverse mask inputs.
+Benefiting from our mask augmentation strategy and Adaptive Target-Aware Attention mechanism, our network demonstrates strong robustness to diverse mask inputs.
 In this subsection, we simulate user strokes and feed them into the network along with the images.
 The resulting outputs and attention maps show that our network can accurately identify and attend to the object and its associated effects, even with imprecise masks, as illustrated in Fig.
-13.
+15.
 Figure 15: Results with User Strokes.
 We simulate user strokes and feed them into the network along with the images.
-The resulting outputs and attention maps show that our network can accurately identify and attend to the object and its associated effects, even with imprecise masks.
-C.4 More Comparisons on In-the-wild Data
-In this subsection, we compare ObjectClear with state-of-the-art methods, including SDXL-INP [1], PowerPaint [43], Attentive Eraser [31], DesignEdit [10], and RORem [15], on in-the-wild data.
-To ensure a fair comparison, we consider two settings: (1) All methods use the same object mask; (2) While other methods use finely annotated object-effect masks, our ObjectClear still uses the object mask.
-The results under these two settings are shown in Fig.
-14 and Fig.
-15 separately.
-Under the first setting, when given the object mask only, our ObjectClear effectively removes shadow and reflection associated with the target object, while all the other methods fail to remove those effects and tend to hallucinate undesirable content or create artifacts.
-Under the second setting, while other methods do remove the shadow thanks to the given effect region, they also alter or remove the original background content in an undesirable way.
-In contrast, our ObjectClear removes the object effect accurately with the object mask only while preserving the background with high fidelity.
-Figure 14: Object Removal on OBER-Wild - Condition on Object Mask.
-Given the object mask only, our ObjectClear effectively removes shadow and reflection associated with the target object, while all the other methods fail to remove those effects and tend to hallucinate undesirable content or create artifacts.
-Figure 15: Object Removal on OBER-Wild - Comparison between methods conditioned on Object-Effect Mask (others) and Object Mask (ours).
-Since existing methods struggle to remove object effects when provided with only the object mask, we supply them with our annotated object-effect masks for a fair comparison.
+The resulting outputs and attention maps show that our network can accurately identify and attend to the object and its effects, even with imprecise masks.
+8.7 Object Insertion and Movement.
+As shown in Fig. 16, even when only the target objects are specified for insertion and movement, ObjectClear is capable of generating plausible and natural shadows and reflections accordingly.
+Figure 16: Object Insertion and Movement. In addition to accurately inserting or repositioning objects, our ObjectClear also generates plausible and natural shadows and reflections accordingly.
+8.8 More Comparisons on In-the-wild Data
+In this subsection, we conduct comprehensive comparisons of ObjectClear with state-of-the-art methods across two categories: image inpainting methods and object removal methods, on the in-the-wild data.
+Comparison with Image Inpainting Methods.
+We compare our method with state-of-the-art image inpainting methods, including PowerPaint [48], Attentive Eraser [34], DesignEdit [10], and RORem [18].
+Since these inpainting methods often struggle to remove object-associated effects when only provided with an object mask, we supply them with our annotated Object-Effect Masks for fair comparison, while our ObjectClear uses only Object Masks as input.
+As shown in Fig.
+20, although these inpainting methods can remove shadows with the guidance of additional effect regions, they frequently introduce undesirable changes to the original background.
+In contrast, ObjectClear effectively eliminates the target object and its associated effects using only the object mask, while preserving background content with high fidelity.
+Comparison with Object Removal Methods.
+We further compare ObjectClear with object removal methods (OmniEraser [37] and OmniPaint [42]) under a unified input setting, all methods use Object Masks as input to ensure fairness.
+As illustrated in Fig.
+19, our ObjectClear demonstrates superior stability compared to the two baselines: it not only successfully eliminates the target object and its associated effects but also avoids unintended changes to the original background, preserving background fidelity with high precision.
+To further highlight the background preservation advantage, we visualize the pixel-wise difference maps of the compared methods in Fig.
+20.
+ObjectClear maintains minimal pixel differences in non-target background areas, which clearly validates its ability to preserve the original background while accurately removing the object and its effects.
+Efficiency Comparison with Object Removal Methods.
+ObjectClear is significantly more efficient than both OmniEraser [37] and OmniPaint [42].
+While the latter two are built on Flux ($\sim 12$B parameters), ObjectClear is based on SDXL ($\sim 3.7$B), making our model considerably lighter.
+As a result, ObjectClear runs over $5\times$ faster during inference, i.e., 1.63 s/image (ObjectClear) vs.
+9.42 s/image (OmniPaint) at $512\times 512$ resolution on an A100 GPU.
+8.9 Limitations
+While ObjectClear exhibits strong performance in removing objects and their associated effects, it still faces challenges in highly complex scenarios.
+Specifically, in cases with overlapping shadows from multiple objects or complex lighting conditions, it can be difficult to disentangle which shadows belong to which objects.
+As a result, the model may fail to remove the shadows of the target object (Fig.
+17a) or remove shadows of other objects (Fig.
+17b).
+Effectively disentangling object-specific shadows in such complex scenes remains an important direction for future work.
+Figure 17: Limitations.
+In complex scenes where multiple objects may produce overlapping or intertwined effects (e.g., shadows and reflections), our method can consistently remove the objects but may sometimes fail to precisely eliminate the associated effects.
+This results in either (a) under-removal of the target effect or (b) over-removal of effects belonging to nearby objects.
+Figure 18: Comparison with Image Inpainting Models.
+Since some methods struggle to remove object effects when provided with only the object mask, we supply them with our annotated Object-Effect Mask for a fair comparison, while our ObjectClear uses only Object Mask as input.
 Although these methods are able to remove shadows with the additional effect region, they often introduce undesirable changes to the original background.
 In contrast, ObjectClear effectively removes the object and its associated effects using only the object mask, while preserving the background content with high fidelity.
-C.5 Limitations
-While ObjectClear exhibits strong performance in removing objects and their associated effects, it still faces challenges in highly complex scenarios.
-Specifically, (1) when multiple shadows heavily overlap, the model struggles to accurately disentangle and localize the effect corresponding to the target object; and (2) in cases involving extremely large and long shadow regions, ObjectClear may fail to completely remove the effect, occasionally leaving behind residual artifacts.
-Enhancing the disentanglement of complex effect regions and improving the removal of large-scale effects remain important directions for future research.
+Figure 19: Comparison with Object Removal Models.
+All compared methods use Object Masks as input for fair comparison.
+Our ObjectClear demonstrates superior stability: it not only successfully eliminates the target object and its associated effects, but also avoids unintended changes to the original background, preserving background fidelity with high precision.
+Figure 20: Comparison of Object Removal Models via Difference Maps.
+We present results of representative object removal methods, and their pixel-wise difference maps relative to the input.
+ObjectClear successfully removes the target object and its associated shadows while preserving the background.
+This advantage is clearly reflected by the low pixel-wise differences in non-target areas.